### v1

# GeoLangBind: Unifying Earth Observation with Agglomerative Vision–Language Foundation Models

Abstract
Earth observation (EO) data, collected from diverse sensors with varying imaging principles, present significant challenges in creating unified analytical frameworks. We present GeoLangBind, a novel agglomerative vision–language foundation model that bridges the gap between heterogeneous EO data modalities using language as a unifying medium. Our approach aligns different EO data types into a shared language embedding space, enabling seamless integration and complementary feature learning from diverse sensor data. To achieve this, we construct a large-scale multimodal image–text dataset, GeoLangBind-2M, encompassing six data modalities. GeoLangBind leverages this dataset to develop a zero-shot foundation model capable of processing arbitrary numbers of EO data channels as input. Through our designed Modality-aware Knowledge Agglomeration (MaKA) module and progressive multimodal weight merging strategy, we create a powerful agglomerative foundation model that excels in both zero-shot vision–language comprehension and fine-grained visual understanding. Extensive evaluation across 23 datasets covering multiple tasks demonstrates GeoLangBind’s superior performance and versatility in EO applications, offering a robust framework for various environmental monitoring and analysis tasks. The dataset and pretrained models will be publicly available.
1 Introduction
Figure 1: GeoLangBind leverages language as a bridge to unify heterogeneous remote sensing data, enabling a common representation space across diverse modalities for improved multimodal learning.
Earth observation (EO) technologies have undergone remarkable advancement in recent years, generating vast amounts of multimodal geospatial data spanning different sensors and scales [54, 55, 70, 59]. In response to the exponential growth of remote sensing data, researchers have proposed foundation models pretrained on large-scale EO data, including SatMAE [12], Scale-MAE [38], HyperSIGMA [46], SSL4EO [47, 39], and SpectralGPT [20]. However, these models typically focus on a single EO data modality, limiting their ability to leverage complementary characteristics across different sensor types.
To address this limitation, recent works such as OFA-Net [58], DOFA [57], AnySat [5], OmniSat [6], and FoMo-Net [8] have introduced foundation models capable of learning from multiple EO modalities within a unified framework. These approaches have demonstrated promising results in multimodal EO representation learning. However, their focus remains on unsupervised visual pretraining, lacking a structured way to align EO data with language for better interpretability and retrieval. Meanwhile, multimodal large language models like GPT [33], LLaMA [44], LLaVA [29], and DeepSeek [27] have shown remarkable capabilities in handling diverse and complex tasks. Vision-language foundation models, such as CLIP [36] and SigLIP [64], have played a crucial role as vision encoders of such multimodal large language models. Inspired by these advances, EO-specific adaptations like RemoteCLIP [28], RS5M [67], and SkyScript [50] have attempted to bring contrastive learning to remote sensing. However, these models are primarily limited to high-resolution RGB optical imagery, and can not handle other diverse EO modalities (e.g., multispectral data, Synthetic Aperture Radar data) that are critical for many applications.
Training a single vision–language foundation model capable of supporting multiple heterogeneous EO data modalities presents several challenges: 1) the lack of comprehensive multimodal EO datasets with aligned textual descriptions, 2) the design of a flexible architecture to handle variable-length input channels across different modalities, and 3) addressing the imbalance among different modalities in the training data. To address such challenges, we propose GeoLangBind, a unified multimodal vision–language foundation model that handles diverse data modalities and aligns them to the shared language feature space, as illustrated in Fig. 1. In addition, GeoLangBind enhances fine-grained semantic understanding via knowledge agglomeration [37] from multiple teacher models beyond contrastive learning.
To achieve this, we combine the following key designs: 1) a wavelength-aware dynamic encoder to handle multiple data modalities, 2) a modality-aware knowledge agglomeration module, and 3) progressive weight-space merging [53] to scale the model efficiently across multiple modalities. Our contributions are as follows:
  1. 1.
    We construct GeoLangBind-2M, a comprehensive dataset of two million image–text pairs spanning six distinct EO data modalities to facilitate the training of multimodal geospatial foundation models.
  2. 2.
    We propose an effective Modality-aware Knowledge Agglomeration (MaKA) module to enhance continuous pre-training and improve fine-grained image understanding.
  3. 3.
    We design a progressive weight merging strategy that enables scalable model adaptation across multiple modalities, outperforming traditional data mixing approaches.
  4. 4.
    We conduct extensive experiments on zero-shot classification, semantic segmentation, and cross-modal image retrieval, demonstrating the superiority of GeoLangBind across diverse tasks.
Benchmarking results indicate that GeoLangBind achieves state-of-the-art performance while being flexible in handling diverse data modalities with any number of spectral bands.
Figure 2: Visualization of data samples from our GeoLangBind-2M. The dataset includes imagery from six different sensors and modalities: Sentinel-2 multispectral, Sentinel-1 SAR, EnMAP hyperspectral, elevation maps, infrared imagery, and aerial imagery. Each sample is paired with textual descriptions capturing key land cover types, objects, and geographic features.
2 Related work
Vision foundation models in Earth observation
Existing pretrained models predominantly focus on RGB data, such as GFM [31], Scale-MAE [38], and Cross-Scale MAE [43]. Others specialize in multispectral imagery, with FG-MAE [48] and SatMAE [12] tailored for Sentinel-2 data, while SSL4EO-L [39] is trained on Landsat imagery. SpectralGPT [20] leverages a 3D generative transformer for spectral data analysis. Beyond optical data, CROMA [14] introduces dual unimodal encoders for multispectral and Synthetic Aperture Radar (SAR) data, using a cross-modal radar–optical transformer to learn unified deep representations. DeCUR [49] enhances bimodal learning by decoupling unique and shared features across modalities. Meanwhile, Satlas [7] assembles a large-scale multi-sensor dataset, pretraining separate models for each sensor. DOFA [57] introduces a neural plasticity-inspired hypernetwork that adapts to different sensor wavelengths, enabling joint training on five distinct sensors. OmniSat [6] proposes to fuse features across multiple EO modalities to learn robust multimodal representations without labeled data. SkySense [18] utilizes separate visual encoders for different data modalities and designs a dedicated module for multimodal fusion.
Zero-shot foundation models
Zero-shot foundation models like CLIP [36] have significantly advanced computer vision by enabling models to generalize across tasks without explicit task-specific training. Building upon CLIP, SigLIP [64] introduces a pairwise sigmoid loss function to enlarge the training batch and enhance the model performance. Zero-shot foundation models have also been investigated in EO applications. SkyScript [50] constructs a large-scale image–text dataset by linking remote sensing images with OpenStreetMap [34] semantics. RemoteCLIP [28] introduces an image–text dataset by generating captions from existing annotated datasets. GeoRSCLIP [67] introduces RS5M, an image–text dataset with 5M pairs, created by filtering existing datasets and generating captions for label-only remote sensing data. Despite these advances, existing vision–language models mainly focus on RGB data and have limitations in learning representations from multiple EO data modalities, which are crucial for holistic environmental analysis. To unify multiple data modalities, ImageBind [16] aligns six modalities within a shared embedding space. LanguageBind [69] extends this by using language as a universal intermediary for modality alignment. While these approaches demonstrate the potential of language or vision as unifying mediums, applying them to EO data is still challenging due to the scarcity of multimodal image–text datasets and the diverse scales of Earth systems.
3 Dataset construction: GeoLangBind-2M
Table 1: Summary of dataset sources of GeoLangBind-2M. * indicates image–text datasets that are created in this work. SAR = synthetic aperture radar, MSI = multispectral imagery, HSI = hyperspectral imagery, and IR = infrared.
Image–Text Datasets Knowledge Domain # Samples
Seg4 General segmentation 41,172
Det10 General detection 110,800
Flair2-RGB-caption* Land cover analysis 61,711
VRS-train General detection 20,264
Skyscript dataset OpenStreetMap tags 1,518,888
SAR Datasets
MMflood-caption* Flood mapping 6,181
SAR-ship-caption* Ship detection 5,984
ChatEarthNet-SAR-caption* Land cover analysis 95,620
MSI/HSI Datasets
ChatEarthNet-S2-caption* Land cover analysis 95,620
NLCD-hyper-caption* Land cover analysis 15,000
Elevation Dataset
Flair2-Elevation-caption* Digital Elevation Model 61,711
Infrared Dataset
IR-ship-caption* Ships in infrared imagery 18,032
Total 2,050,983
GeoLangBind-2M contains a broad range of data modalities including RGB, SAR, multispectral, hyperspectral, infrared, and elevation data, as listed in Tab. 1. Each image is paired with a high-quality text annotation, as illustrated in Fig. 2. All datasets are implemented using the TorchGeo library [40]. In the following, we introduce the composition of the dataset.
RGB image–text datasets
High-resolution RGB images with textual descriptions form a significant portion of GeoLangBind-2M. We incorporate parts of the RemoteCLIP [28] dataset, i.e. Seg4 and Det10, which are constructed by generating captions from existing segmentation and detection datasets. We also include VRSBench [26], which aggregates multiple object detection datasets. SkyScript [50] is a large-scale image–text dataset for EO, from which we include the top 30% highest quality samples after filtering. Beyond existing datasets, we introduce Flair2-RGB-caption [15], a newly created image–text dataset. Using semantic information as prompts, we employ the Pixtral-12B [2] model to generate captions for detailed land cover features.
SAR image–text datasets
To enhance the dataset with SAR–text pairs, we create MMFlood-SAR-caption [32], which contains captions emphasizing water extent and flooded regions. Additionally, the SAR-ship-caption dataset includes captions describing ships in SAR images. ChatEarthNet-SAR-caption further extends the collection by covering diverse SAR scenes from Sentinel-1, focusing on land cover types.
Multispectral, hyperspectral, and elevation image–text datasets
We construct ChatEarthNet-S2-caption by summarizing long, detailed descriptions from the ChatEarthNet dataset [63]. This dataset contains Sentinel-2 multispectral imagery covering more than 10 land cover types. Furthermore, we introduce NLCD-hyper-caption, a hyperspectral dataset with 200+ EnMAP bands [17]. Like Flair2-RGB-caption, we use Pixtral-12B to generate rich captions, leveraging annotated land cover maps for enhanced descriptiveness. Similarly, we also generate captions using Pixtral-12B to construct the Flair2-Elevation-caption dataset. Further details on the construction of GeoLangBind-2M are provided in the Supplementary materials.
Figure 3: GeoLangBind aims to unify and align diverse Earth observation data modalities into the shared language feature space while enhancing fine-grained image understanding. The model uses modality-aware knowledge agglomeration (MaKA) to transfer knowledge from multiple teachers. The wavelength-aware dynamic patch embedding module processes multimodal images dynamically.
4 Methodology
4.1 GeoLangBind model design
GeoLangBind aims to unify diverse data modalities and embed them into a shared language feature space while enhancing fine-grained image understanding ability. To this end, we design a wavelength-aware dynamic encoder, a modality-aware knowledge agglomeration module, and a progressive weight-space merging approach.
4.2 Wavelength-aware dynamic encoder
To handle the diversity of spectral bands across modalities, we adopt the dynamic encoder architecture from DOFA [57], as illustrated in Fig. 3. Given an input image $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral channels, we define its corresponding central wavelengths as $\mathbf{\lambda}\in\mathbb{R}^{C}$. For modalities without well-defined wavelengths, such as elevation, we assign default wavelengths equivalent to those of the RGB channels. First, a 1D sine-cosine positional encoding is applied to embed these wavelengths into a higher-dimensional space: $\mathbf{V}_{\lambda}=\text{PE}(\mathbf{\lambda})\in\mathbb{R}^{C\times d_{\lambda}},$ where $d_{\lambda}$ is the embedding dimension. The encoded wavelengths $\mathbf{V}_{\lambda}$ are then transformed using two fully-connected layers with residual connections.
Next, a lightweight Transformer [45] encoder processes the transformed wavelength embeddings. Specifically, we concatenate the transformed embeddings $\mathbf{V^{\prime}}_{\lambda}$ with the learnable weight query tokens $\mathbf{Q}_{w}$ and a learnable bias query token $\mathbf{Q}_{b}$ as input to the Transformer encoder. After the Transformer layer, the output features are passed through two fully-connected layers to generate the dynamic convolutional weights $\mathbf{M}_{w}$ and biases $\mathbf{M}_{b}$. Finally, these dynamically generated weights and biases are used for patch embedding (implemented as convolution). This process enables the model to handle different data modalities with varying spectral channels using a unified deep encoder.
4.3 Modality-aware knowledge agglomeration
We introduce the modality-aware knowledge agglomeration (MaKA) module, which leverages wavelength as a modality-specific condition when distilling features from different data modalities. MaKA comprises three key components: a wavelength-aware prompt generator, which encodes wavelength values into a learnable embedding vector; a modality-aware conditional layer normalization, which adapts normalization parameters based on modality-specific characteristics; and a feature projector, which transforms input features to facilitate effective feature matching and knowledge distillation. Below, we introduce each of them in detail.
Wavelength-aware prompt generator
Each modality is associated with a set of wavelengths, which we encode using a sine-cosine positional encoding function. Given input wavelengths $\mathbf{\lambda}\in\mathbb{R}^{C}$, we first obtain their positional embeddings $\mathbf{V}_{\lambda}\in\mathbb{R}^{C\times d_{\lambda}}$. To generate the modality-aware prompt vector, we apply a linear projection $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times d_{\lambda}}$ to each row of $\mathbf{V}_{\lambda}$, then compute the mean over the $C$ wavelengths:
$\mathbf{V}_{p}=\frac{1}{C}\sum_{i=1}^{C}\bigl{(}\mathbf{W}_{\text{proj}}\,\mathbf{V}_{\lambda_{i}}\bigr{)},$ (1)
where $\mathbf{V}_{p}\in\mathbb{R}^{d}$ is the fused modality embedding. This final embedding vector $\mathbf{V}_{p}$ captures modality-specific information derived from the provided wavelengths.
Modality-aware layer normalization
Given the feature map from the last layer of the GeoLangBind model, we first interpolate it to match the teacher’s feature resolution, denoted as $\mathbf{F}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$, omitting the batch dimension for simplicity. With $\mathbf{F}$ and a modality prompt $\mathbf{V}_{p}\in\mathbb{R}^{d}$, our goal is to perform layer normalization that dynamically adapts to the modality. To achieve this, we introduce modality-aware conditional layer normalization, formulated as follows.
First, we derive two learnable modulation vectors, $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{d}$, from the modality prompt via a linear transformation:
$\begin{bmatrix}\boldsymbol{\gamma}\\[6.0pt] \boldsymbol{\beta}\end{bmatrix}\;=\;\mathbf{W}_{\text{prompt}}\,\mathbf{V}_{p}\;\in\;\mathbb{R}^{2d},$ (2)
where $\mathbf{W}_{\text{prompt}}\in\mathbb{R}^{2d\times d}$. We then split the output into $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{d}$ and reshape them to broadcast over the spatial dimensions:
[…]
Let $\mathrm{LN}(\mathbf{F})$ denote the standard channel-wise layer normalization of $\mathbf{F}$. The modality-aware layer normalization is then defined as:
$\mathbf{F}^{\prime}\;=\;\mathrm{LN}(\mathbf{F})+\boldsymbol{\gamma}\odot\mathrm{LN}(\mathbf{F})\;+\;\boldsymbol{\beta},$ (4)
where $\odot$ denotes the Hadamard product. Note that we use the modality-aware layer normalization to learn the residual information specific to each modality.
Projector module
A convolutional layer is applied to adjust the feature representation $\mathbf{F}^{\prime}$ to channel dimensions $d_{t}$ of the corresponding teachers. The final representation is then used to compute the feature-matching loss for model distillation. Our design incorporates modality awareness into the feature translation process by encoding wavelength information and adapting the layer normalization parameters accordingly. This ensures that the learned features are both aligned and modality-specific, enhancing robust adaptation across different input modalities.
Model training
To train the model, we use a feature-matching loss to align the translated spatial features with the teacher’s features. As illustrated in Fig. 3, we extract three RGB channels from multispectral or hyperspectral images as input to the teacher model. The student features $\mathbf{F}_{s}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ are then matched to the teacher’s features $\mathbf{F}_{t}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ using a combination of $L_{1}$ loss, cosine embedding loss to maximize the cosine similarity, and mean squared error (MSE) loss:
$L_{\text{match}}=L_{L_{1}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{mse}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{cos}}(\mathbf{F}_{s},\mathbf{F}_{t}).$ (5)
We utilize three pre-trained models as teachers: SigLIP [64], DINOv2 [35], and ViT [13]. During training, the GeoLangBind model is initialized with SigLIP pre-trained weights and trained on GeoLangBind-2M using both the pairwise sigmoid loss $L_{\text{siglip}}$ for contrastive learning and distillation loss for feature matching. The total loss function is:
$L=L_{\text{siglip}}+\alpha_{s}L_{\text{match}}^{\text{siglip}}+\alpha_{d}L_{\text{match}}^{\text{dinov2}}+\alpha_{v}L_{\text{match}}^{\text{vit}},$ (6)
where $\alpha_{s},\alpha_{d},$ and $\alpha_{v}$ are balancing factors for feature-matching losses of SigLIP, DINOv2, and ViT, respectively.
Table 2: Zero-shot comparison of various models on scene and fine-grained classification tasks. Bold values indicate the best performance.
ViT Model Scene classification Fine-grained classification
SkyScript AID EuroSAT fMoW Million-AID PatternNet RESISC RSI-CB Avg. Roof shape Smoothness Surface
Base CLIP-original 40.16 69.55 32.11 17.62 57.27 64.09 65.71 41.26 49.66 31.50 26.80 61.36
Human-curated captions 40.03 71.05 33.85 18.02 57.48 66.56 66.04 42.73 50.82 28.50 27.80 60.91
RemoteCLIP 27.06 87.05 30.74 11.13 46.26 56.05 67.88 44.55 49.09 30.50 21.00 43.86
CLIP-laion-RS 40.77 69.55 37.63 19.16 56.59 64.79 64.63 41.79 50.59 28.83 27.60 62.27
SkyCLIP-50 52.98 70.90 33.30 19.24 62.69 72.18 66.67 46.20 53.02 26.00 38.00 67.73
[…]
Large CLIP-original 55.06 69.25 41.89 26.19 57.88 71.39 66.70 43.02 53.76 37.50 25.40 42.73
Human-curated captions 56.09 72.95 41.96 26.33 58.47 74.86 68.70 44.60 55.41 37.00 26.60 40.00
RemoteCLIP 34.40 70.85 27.81 16.77 47.20 61.91 74.31 50.79 49.99 34.33 34.20 55.45
CLIP-laion-RS 58.81 71.70 54.30 27.21 60.77 72.68 71.21 48.21 57.82 40.50 37.60 53.41
SkyCLIP-20 67.94 71.95 53.63 28.04 65.68 78.62 70.70 50.03 59.98 44.83 26.80 61.36
SkyCLIP-30 69.08 72.15 52.44 27.77 66.40 79.67 70.77 50.19 59.99 46.17 30.80 64.32
SkyCLIP-50 70.89 71.70 51.33 27.12 67.45 80.88 70.94 50.09 59.93 46.83 35.80 67.50
[…]
Table 3: Zero-shot comparison of various models on GeoBench datasets. Bold values indicate the best performance.
Model (Base version) m-bigearthnet m-so2sat m-forestnet
Precision Recall F1 RGB Sentinel-2 (5 bands) RGB Landsat-8 (5 bands)
SigLIP 45.34 12.36 16.82 12.88 - 8.16 -
RemoteCLIP 33.82 20.35 18.84 10.75 - 8.46 -
SkyCLIP-50 39.91 19.58 20.32 11.97 - 10.78 -
[…]
4.4 Progressive Multimodal Weight Merging
To address the data imbalance issue, we propose a post-training weight-space merging strategy to scale the model to diverse data modalities effectively. Since RGB data is the most dominant modality, we split the dataset into two parts: data with RGB channels and data with other modalities. During training, we first train two separate GeoLangBind models on these subsets. After training, we merge their weights to obtain a unified model capable of handling both RGB and other modalities. Since the original SigLIP model has been trained on massive data, we also merge the weights of GeoLangBind with the original SigLIP weights to enhance the model. Thus, we perform two rounds of weight merging to merge these three model weights: 1) merge the original SigLIP weights with the model trained on the RGB subset of GeoLangBind-2M (GeoLangBind-RGB); 2) merge the intermediate model with the model trained on the remaining five data modalities (GeoLangBind-Others).
For clarity, let $\theta_{\text{siglip}}$ be the original SigLIP model weights, $\theta_{\text{rgb}}$ be the GeoLangBind model weights trained on the RGB subset, $\theta_{\text{others}}$ be the GeoLangBind model weights trained on the other five modalities. In the first step, we merge $\theta_{\text{siglip}}$ and $\theta_{\text{rgb}}$ using a simple linear weight merging strategy:
$\theta^{*}=(1-m_{1})\,\theta_{\text{siglip}}+m_{1}\,\theta_{\text{rgb}},$ (7)
where $m_{1}$ controls the weighting ratio between the two models. Next, we merge the intermediate weights $\theta^{*}$ with $\theta_{\text{others}}$ to obtain the final GeoLangBind model:
$\theta=(1-m_{2})\,\theta^{*}+m_{2}\,\theta_{\text{others}},$ (8)
where $m_{2}$ determines the contribution of the non-RGB modalities. We conduct ablation studies in Sec. 5.3 by varying $m_{1}$ and $m_{2}$ to search for optimal weighting ratios. Although more sophisticated weight merging methods exist [3], we adopt a simple linear approach to evaluate its effectiveness for heterogeneous EO data. Investigating advanced merging strategies remains a topic for future work.
5 Experiments
5.1 Implementation details
GeoLangBind is trained using 8× NVIDIA A100 40GB GPUs for the base model and 8× NVIDIA A100 80GB GPUs for the large version. Our code is implemented based on OpenCLIP [21]. For the base version, we use a ViT-Base architecture (patch size 16), initialized with WebLI-pretrained weights [9]. The teacher models are ViT-Base-SigLIP (224), DINOv2-ViT-L/14, and ViT-Huge (ImageNet). The large version is based on SoViT-400m[4] (patch size 14, input size 384), with teacher models ViT-So400m-SigLIP (384), DINOv2-ViT-L/14, and ViT-Large (ImageNet). Training is conducted for 5 epochs on RGB datasets and 20 epochs on other modalities using AdamW (learning rate 5e-4, weight decay 1e-7). The batch size is 140 for the base model and 35 for the large model.
5.2 Performance comparison
We follow the evaluation protocols of SkyScript [50] and evaluate the models via zero-shot classification tasks. Specifically, the following eleven image scene classification datasets are used: AID [54], EuroSAT [19], fMoW [11], Million-AID [30], PatternNet [68], NWPU-RESISC45 [10], RSI-CB256 [25], as well as the fine-grained attribute classification datasets. In addition, we further compare these models on some datasets from the GEO-Bench suite [24]: m-bigearthnet [41], m-so2sat [71], and m-forestnet [22].
Table 4: Recall@1, recall@5, and recall@10 for cross-modal retrieval on three benchmark datasets. Bold values indicate the best performance.
RSICD RSITMD UCM-caption
Image to Text Text to Image Image to Text Text to Image Image to Text Text to Image
Model R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
RSICD, RSITMD, and/or UCM-caption seen in training
AMFMN [61] 5.39 15.08 23.40 4.90 18.28 31.44 10.63 24.78 41.81 11.51 34.69 54.87 16.67 45.71 68.57 12.86 53.24 79.43
LW-MCR-u [60] 4.39 13.35 20.29 4.30 18.85 32.34 9.73 26.77 37.61 9.25 34.07 54.03 18.10 47.14 63.81 13.14 50.38 79.52
GaLR [62] 6.59 19.85 31.04 4.69 19.48 32.13 14.82 31.64 42.48 11.15 36.68 51.68
RSICD, RSITMD, and/or UCM-caption not seen in training
CLIP-original 6.59 20.68 31.75 3.62 14.28 23.63 10.18 30.31 42.04 8.31 24.96 39.03 37.62 78.10 89.52 28.12 64.99 77.19
CLIP-laion-RS 8.42 23.70 35.86 5.81 19.49 30.25 13.72 32.08 44.91 10.57 31.48 46.96 39.52 79.52 90.00 29.71 62.60 80.37
SkyCLIP-30 8.97 24.15 37.97 5.85 20.53 33.53 11.73 33.19 47.35 10.19 32.47 49.08 38.57 84.29 93.81 31.83 64.19 81.96
[…]
Table 5: Partial fine-tuning results on six segmentation tasks. All models are trained with a frozen backbone for 20 epochs. Reported numbers are mean intersection over union (mIoU). Missing values are due to the inability of the model to adapt to this domain.
Method
Models
Backbone
m-pv4ger-seg
m-nz-cattle
m-NeonTree
m-cashew-plant
m-SA-crop
m-chesapeake
Supervised DeepLabv3 ResNet101 93.4 67.6 53.9 48.6 30.4 62.1
U-Net ResNet101 94.1 80.5 56.6 46.6 29.9 70.8
Self-supervised Vision FMs rand. init. ViT-B 81.7 74.1 51.7 32.4 29.0 47.1
Scale-MAE [38] ViT-L 83.5 76.5 51.0 - - 61.0
GFM [31] Swin-B 92.0 75.0 51.1 - - 63.8
Cross-Scale MAE [43] ViT-B 83.2 77.9 52.1 - - 52.3
CROMA [14] ViT-B - - - 30.1 31.4 -
DOFA [57] ViT-B 94.5 81.4 58.8 51.5 33.0 65.3
DOFA [57] ViT-L 95.0 81.8 59.4 56.9 32.1 66.3
Zero-shot FMs RemoteCLIP [28] ViT-L 94.1 79.8 56.2 37.9 31.3 59.5
Skyscript-20 [50] ViT-L 93.9 79.2 57.2 38.0 30.6 61.5
GeoLangBind-B-224 ViT-B 94.3 80.7 56.1 53.6 31.8 62.7
GeoLangBind-L-384 Large 94.7 82.2 59.0 45.8 32.6 63.0
Overall zero-shot performance
As shown in Tab. 2, our GeoLangBind models establish new state-of-the-art results in zero-shot remote-sensing image classification. Compared with existing CLIP-based baselines such as CLIP-original, RemoteCLIP, CLIP-laion-RS, and the various SkyCLIP configurations, GeoLangBind-B-224 (Base) consistently outperforms all counterparts on scene-level tasks. Its average accuracy across the eight scene datasets is significantly higher than that of the other Base models. The Larger variant of GeoLangBind with higher resolution yields further accuracy gains. This highlights that both capacity and resolution play vital roles in capturing the spatial and spectral details crucial for complex remote-sensing tasks. On three fine-grained roof-attribute classification tasks, both GeoLangBind variants achieve strong accuracy on roof shape and surface, significantly outperforming other methods. Although the results on smoothness remain more modest, our approach still provides a clear improvement over existing approaches.
Zero-shot classification on GEO-Bench
The zero-shot classification results on three GEO-Bench datasets are presented in Tab. 3. GeoLangBind demonstrates competitive or superior zero-shot multi-label classification performance compared to existing models on the m-bigearthnet dataset (43 classes). For the m-so2sat dataset (17 classes), when utilizing five spectral bands, the performance decreases slightly due to the significantly less multispectral training data compared to RGB. Note that competitive models cannot process 5-band images as input for zero-shot classification. On the m-forestnet dataset (12 classes), GeoLangBind substantially outperforms other models and shows remarkable accuracy improvements when using five spectral bands as input. Notably, the m-forestnet dataset uses Landsat-8 satellite data, which is not represented in our training dataset.
Cross-modal retrieval performance
Tab. 4 compares the GeoLangBind model against recent methods on three benchmark caption datasets: RSICD, RSITMD, and UCM-caption. We report Recall@1, Recall@5, and Recall@10 for both image-to-text and text-to-image retrieval. Despite the inherent difficulty of remote-sensing cross-modal tasks, GeoLangBind-L consistently achieves competitive or higher recall scores than most competitive baselines, including CLIP-original, CLIP-laion-RS, and SkyCLIP-30.
Semantic segmentation performance
Tab. 5 shows that GeoLangBind models work well across six segmentation tasks despite being zero-shot vision–language models. For all models, the backbone remains frozen while UPerNet [56] is trained as the segmentation head. GeoLangBind-B-224 achieves strong mIoU scores on m-pv4ger-seg and m-nz-cattle datasets. GeoLangBind-L-384 further improves results across all datasets, achieving competitive or better performance than vision foundation models. These results demonstrate GeoLangBind’s versatility and ability on fine-grained pixel-level understanding tasks.
Table 6: Ablation experiments of models on the AID zero-shot classification dataset.
Model Distill. Top-1 Accuracy Top-5 Accuracy
SigLIP 65.95 96.20
[…]
SigLIP+GeoLangBind-B-RGB 77.70 95.50
SigLIP+GeoLangBind-B-RGB 72.00 95.80
[…]
SigLIP+GeoLangBind-B-Others 69.50 96.20
SigLIP+GeoLangBind-B-Others 66.30 91.20
[…]
GeoLangBind-B-All-Merging 77.60 96.65
5.3 Ablation studies
Table 7: Ablation experiments of different distillation methods on the segmentation datasets.
Model MADOS m-nz-cattle m-NeonTree
SigLIP 57.2 77.5 52.3
[…]
GeoLangBind-B w/ MaKA 62.3 82.2 59.0
Table 8: Ratio for merging SigLIP and GeoLangBind-B-RGB.
$m_{1}$ AID (Top-1) RSI-CB (Top-1) Roof shape (Top-1)
0 65.95 38.01 41.17
0.1 44.25 27.33 37.67
0.3 56.75 31.87 38.00
0.5 74.65 41.80 44.33
0.7 77.40 45.36 53.67
0.9 77.70 47.03 57.83
1.0 76.90 46.72 58.50
Weight merging
Tab. 6 presents the results of our ablation experiments, evaluating the impact of weight merging across different models on the AID zero-shot classification dataset. The baseline SigLIP model can be significantly improved by incorporating the GeoLangBind-B-RGB model, trained on the RGB part of GeoLangBind-2M. Merging SigLIP with GeoLangBind-B-RGB further enhances performance. The model trained with our MaKA module (for distillation) can largely outperform the non-distilled version. This highlights the role of distillation in improving weight fusion. GeoLangBind-B-Others, trained on non-RGB modalities performs lower than its RGB counterpart. When merged with SigLIP, the performance can be significantly improved.
GeoLangBind-B-All is the model trained with all the data modalities mixing together, i.e. using data mix instead of weight merging. When compared with the GeoLangBind-B-All-Merging, the performance is significantly lower. This comparison demonstrates that weight merging is more effective and flexible than direct mixing data modalities.
Knowledge distillation
As presented in Tab. 7, we conducted ablation studies on the MADOS [23], m-nz-cattle [1], and m-NeonTree [52] datasets. GeoLangBind-B without MaKA represents a model trained using a baseline distillation method that does not incorporate modality-specific wavelengths as conditioning. The results demonstrate the significant effectiveness of the MaKA module, showing consistent performance improvements across all datasets.
Weight merging ratios
In Tab. 8 and Tab. 9, a linear search is conducted to determine the optimal ratios for weight merging. Based on the results, we select 0.9 for $m_{1}$ to merge the weights of SigLIP and GeoLangBind-RGB. We choose 0.5 for merging weights of the intermediate weights and GeoLangBind-Others. Thus we set $m_{1}=0.9$ and $m_{2}=0.5$ to derive our final model weights. The findings demonstrate that merged weights consistently outperform individual models (when the ratio equals 0 or 1), validating the effectiveness of our weight-merging approach.
Table 9: Ratio for merging weights of GeoLangBind-B-RGB and GeoLangBind-B-Others.
$m_{2}$ AID (Top-1) RSI-CB (Top-1) Roof shape (Top-1)
0 77.70 47.03 57.83
0.1 78.10 47.31 57.50
0.3 78.55 47.91 55.83
0.5 78.60 49.17 52.67
0.7 76.45 49.92 52.00
0.9 76.35 43.98 63.33
1.0 69.20 34.80 59.67
Figure 4: t-SNE visualization of feature distributions for the original SigLIP features (left) and GeoLangBind-L-384 model (right). Circle markers represent language features, while triangle markers represent image features.
Figure 5: Visual comparison of deep features from RemoteCLIP, Skyscript, and GeoLangBind model.
5.4 Feature visualization and analysis
In the left side of Fig. 4, we present the t-SNE plot of SigLIP features. There is a significant separation between the language and image features for hyperspectral, SAR, elevation, and multispectral data modalities. This means that SigLIP struggles to align the representations of these data modalities. In contrast, features of GeoLangBind-L-384 demonstrate a much stronger feature alignment across all modalities, making it a powerful foundation for Earth observation tasks. In Fig. 5, we visually compare the feature heatmaps of RemoteCLIP, SkyScript, and GeoLangBind. We can see that GeoLangBind features are more spatially structured and preserve more object details.
6 Conclusion
We introduce GeoLangBind, a unified vision–language foundation model that bridges heterogeneous EO modalities through a shared language space. To achieve this, we construct GeoLangBind-2M, a large-scale image–text dataset containing six EO data modalities. GeoLangBind consists of a wavelength-aware encoder and a modality-aware knowledge agglomeration module to enhance fine-grained image understanding. Additionally, progressive weight merging is proposed to scale the model training to multiple data modalities. Extensive experiments demonstrate the state-of-the-art performance of GeoLangBind across zero-shot classification, semantic segmentation, and cross-modal retrieval tasks.
Appendix A Supplementary material
The supplementary material contains the following sections:
  1. 1.
    Details of the evaluation datasets;
  2. 2.
    Details on the construction of GeoLangBind-2M;
  3. 3.
    More visualization examples of GeoLangBind-2M;
  4. 4.
    Ablation studies on the distillation loss terms;
  5. 5.
    Confusion matrices of the zero-shot classification tasks;
  6. 6.
    More visualization of the features;
A.1 Details of the evaluation datasets
We evaluate GeoLangBind on both datasets that have been established in previous works for direct comparison, as well as additional datasets that highlight the wide-ranging capabilities of the proposed model.
The following datasets are part of the SkyScript [50] evaluation framework, which we follow for direct comparison.
  • SkyScript classification dataset [50]: contains 7,000 aerial RGB images with 70 classes and different objects than the SkyScript pretraining dataset.
  • AID [54]: contains 2,000 aerial RGB images and 30 image scene classes.
  • EuroSAT [19]: contains 27,000 Sentinel-2 images with 13 spectral bands and 10 image scene classes. We use the dedicated test split that contains 2,700 images.
  • fMoW [11]: contains 106,081 RGB aerial images with 62 image scene classes.
  • Million-AID [30]: contains 10,000 RGB aerial images of varying GSD with 51 image scene classes.
  • PatternNet [68]: contains 30,400 high-resolution RGB aerial images with 6–50 cm GSD and 38 image scene classes.
  • NWPU-RESISC45 [10]: contains 31,500 aerial RGB images with 0.2–30 m GSD and 45 image scene classes.
  • RSI-CB [25]: contains 24,747 aerial RGB images and 35 image scene classes.
Table 10: GEO-Bench dataset description.
Tasks Name Image Size # Classes Train Val Test # Bands Sensors
Classification m-bigearthnet 120 x 120 43 20000 1000 1000 12 Sentinel-2
m-so2sat 32 x 32 17 19992 986 986 18 Sentinel-2 + Sentinel-1
m-forestnet 332 x 332 12 6464 989 993 6 Landsat-8
Segmentation m-pv4ger-seg 320 x 320 2 3000 403 403 3 RGB
m-chesapeake-landcover 256 x 256 7 3000 1000 1000 4 RGBN
m-cashew-plantation 256 x 256 7 1350 400 50 13 Sentinel-2
m-SA-crop-type 256 x 256 10 3000 1000 1000 13 Sentinel-2
m-nz-cattle 500 x 500 2 524 66 65 3 RGB
m-NeonTree 400 x 400 2 270 94 93 5 RGB + Hyperspectral + Elevation
Additionally, we choose the GEO-Bench [24] suite of datasets that cover a range of relevant remote sensing tasks across different domains, sensors, and geospatial locations. Tab. 10 contains a detailed overview of the number of samples, sensors, and target classes in the GEO-Bench datasets.
Figure 6: Word cloud for the Flair2-RGB-caption dataset.
Figure 7: Word histogram for the Flair2-RGB-caption dataset.
In addition, we conduct ablation experiments on the Marine Debris and Oil Spill (MADOS) [23] dataset. The MADOS dataset is a high-resolution multispectral Sentinel-2 dataset for marine pollution detection. Spanning 174 globally distributed scenes (2015–2022) with 1.5M annotated pixels, it covers diverse pollutants, sea surface features, and water-related classes. Unlike existing datasets, MADOS enables scalable, generalizable deep learning models for holistic marine pollution monitoring.
A.2 Details on constructing GeoLangBind-2M
We construct our multimodal image–text dataset GeoLangBind-2M by assembling existing datasets and curating 8 new datasets. Details on the newly curated sub-datasets are as follows.
A.2.1 Flair2-RGB-caption
For the Flair2-RGB-caption dataset, originally designed for semantic segmentation land cover classification, we employed an approach similar to the ChatEarthNet [63] dataset construction. We select the training set of 61,711 images from FLAIR #2 [15] and derive contextual information from their corresponding pixel-level segmentation masks. This extracted semantic content, paired with the segmentation maps as visual context, serves as the input for generating comprehensive descriptions using the Pixtral 12B [2] language model.
Our semantic extraction process involves calculating the distribution percentages of various terrain categories within each image and formulating these statistics into structured contextual cues. We identify all land cover categories in each segmentation map, assigning distinctive color codes to represent different semantic classes. The designed prompt explicitly defines the correspondence between color codes and land cover categories, while also incorporating the calculated percentage coverage of each land/object type within the segmentation maps. This quantitative spatial information is then processed by the Pixtral 12B model to analyze distribution patterns and generate linguistically accurate descriptions.
With this semantic framework established, we craft a prompt for generating detailed image descriptions. The prompt template variables indicated in brackets represent dynamic content populated during the semantic processing phase. The exact prompt template used for Flair2-RGB-caption creation is shown below:
{mdframed}
[backgroundcolor=gray!5, linecolor=black] You are an AI visual assistant capable of describing a scene based on a segmentation map. The map represents different land cover types using specific colors. The legend is as follows:
- [$color\_i$] corresponds to [$label\_i$], occupying [$percent\_i$]of the image.
Do not mention colors, color coding, or technical details. Use the given land cover class names exclusively. Generate a brief yet natural description of the scene by extending the sentence: “The aerial image contains [$presented\_labels$] land types.” Provide a concise summary of their spatial distribution.
In Fig. 6, we illustrate the word cloud for this dataset.
A.2.2 Flair2-Elevation-caption
The Flair2-Elevation-Caption dataset is constructed to generate detailed textual descriptions of terrain characteristics by leveraging both semantic segmentation maps and elevation maps. The dataset creation follows a structured approach, similar to the Flair2-RGB-caption dataset, ensuring that the generated captions effectively describe the topographical and land cover features of each image.
Figure 8: Word cloud for the Flair2-Elevation-caption dataset.
Figure 9: Histogram for the Flair2-elevation-caption dataset. Similarly to the other datasets of this appendix, FLAIR2-Elevation-caption has a focused vocabulary with many rare domain-specific terms. Its distribution deviates from a true power law due to frequent word repetition and an unusually long tail that is populated with remote sensing imagery terms. This mix of common and rare words could help models generalize by balancing domain-specific terms with diverse scene descriptors.
To begin, we select 61,711 images from the FLAIR #2 dataset, each accompanied by a segmentation map and an elevation map. The segmentation map provides land cover information, where each pixel corresponds to a specific land class, while the elevation map encodes height values representing terrain variation. Additionally, a class label mapping is used to associate numerical class IDs with their corresponding semantic labels, such as “forest,” “urban area,” or “water body.” This combination of datasets enables a comprehensive understanding of both land cover distribution and elevation patterns.
For each selected image, we analyze the elevation distribution by identifying the highest and lowest elevation values within the image patch. The corresponding land cover types at these extreme points are then determined using the segmentation map. This extraction process helps characterize the relationship between land cover and elevation, capturing how different landscapes correspond to varying altitude levels.
Using the extracted elevation and land cover information, we design a structured prompt to generate descriptive and natural textual captions. The prompt explicitly states the highest and lowest elevation values and their associated land cover types, followed by a request to summarize the overall terrain characteristics. The model is guided to describe whether the landscape is relatively flat or exhibits significant elevation changes, ensuring that the generated captions provide useful insights for downstream geospatial tasks.
To generate the final dataset, the structured prompts are input into a language model Pixtral 12B, which produces detailed, human-like descriptions of the elevation characteristics. These generated captions are then stored alongside their corresponding elevation and segmentation maps, forming the Flair2-Elevation-Caption dataset. The dataset offers rich, multimodal learning signals, allowing models to develop a deeper understanding of both land cover semantics and terrain variations.
{mdframed}
[backgroundcolor=gray!5, linecolor=black] This is an elevation map that indicates the height of each pixel. The highest areas, at an elevation of approximately [${highest\_height}$] meters, are [$highest\_class$]. The lowest areas, at an elevation of approximately [${lowest\_height}$] meters, are [$lowest\_class$].
Based on the provided context and elevation values, generate a concise and accurate description of the elevation map. Describe the image by briefly addressing:
1) The highest and lowest land cover types.
2) Whether the terrain is relatively flat or has significant elevation differences. In Fig. 8, we demonstrate the word cloud to provide a better overview of the dataset. The histogram of the Flair2-Elevation-caption dataset is presented in Fig. 9.
A.2.3 MMflood-SAR-caption
The MMflood-SAR-caption dataset contains textual descriptions of flood maps generated using Pixtral 12B. In the following, we will describe the generation process in detail. Each flood map is represented as a binary mask, where white pixels indicate flooded areas, and black pixels represent non-flooded regions. To create meaningful textual descriptions, we analyze the extent of flooding and its spatial distribution before constructing structured prompts to guide the language model.
Figure 10: Word cloud for the MMflood-SAR-caption dataset.
Figure 11: Word histogram for the MMflood-SAR-caption dataset.
To quantify the severity of flooding, we calculate the total number of flooded pixels in each binary mask and express it as a percentage of the total image area. This value provides a direct measure of the flood extent, allowing the generated captions to capture variations in flood severity across different samples. In addition to flood coverage, we assess spatial patterns by dividing the image into four quadrants: top, bottom, left, and right. By checking the presence of flooded pixels in each region, we determine whether the flooding is concentrated in specific areas or dispersed across the map.
Once these characteristics are extracted, we construct a structured prompt to guide Pixtral 12B in generating concise and informative descriptions. The prompt instructs the model to summarize the proportion of flooded areas and specify their locations within the image. To ensure clarity and brevity, the generated description is limited to fewer than 70 words, making it suitable for vision–language pretraining.
The prompts are then fed into Pixtral 12B, which generates natural language captions describing each flood map. These captions, stored alongside their corresponding binary masks, form the MMFlood dataset, enabling automated flood analysis, disaster monitoring, and geospatial AI research. By combining semantic flood mapping with natural language descriptions, this dataset enhances the ability of AI models to interpret and describe flood-affected regions in a human-readable format. We also provide the word cloud in Fig. 10 to show the detailed content of this dataset. In Fig. 11, we present the histogram of the words in the MMflood-SAR-caption dataset.
{mdframed}
[backgroundcolor=gray!5, linecolor=black] This is a flood map where areas marked with white pixels indicate flooded regions. The flooded area occupies approximately [$flood\_percentage$]of the entire map. Please analyze the flood map and provide insights into the affected areas.
You must generate a short description (less than 70 words) of the elevation image. First, describe the portion of floods; then introduce the location of the flooded areas.
After generating these detailed textual descriptions, we further refine the dataset by prompting Pixtral 12B to summarize each description into a concise 30-word caption. The prompts used in this summarizing process are as follows:
{mdframed}
[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
A.2.4 NLCD-hyper-caption
The caption generation process for hyperspectral images follows a methodology similar to the FLAIR #2 dataset, ensuring consistency in semantic information extraction and description generation. Hyperspectral images provide rich spectral information across multiple wavelengths, allowing for a more detailed classification of land cover types. However, the caption generation process remains aligned with the Flair2-RGB-Caption dataset, differing only in the specific prompt format used to guide the Pixtral 12B model.
Figure 12: Word cloud for the NLCD-hyper-caption dataset.
Figure 13: Word histogram for the NLCD-hyper-caption dataset.
To create captions, we first extract semantic information from the hyperspectral images using their corresponding segmentation maps. Each land cover type is identified and labeled, and its spatial distribution within the image is analyzed. Unlike conventional RGB imagery, where color-based segmentation might play a role, hyperspectral images rely on spectral signatures to differentiate land cover types, making this process crucial for accurate caption generation.
Once the land cover labels and their distributions are determined, we construct a structured prompt to generate natural language descriptions. The Pixtral 12B model is tasked with producing concise and human-readable descriptions of the scene. The prompt ensures that the generated captions focus solely on land cover types and their spatial distributions, avoiding any mention of technical details, color coding, or hyperspectral data intricacies.
Following the generation of detailed textual descriptions, we further prompt Pixtral 12B to refine the descriptions into brief and natural captions. This step ensures that the dataset provides both long-form and short-form textual annotations, improving its usability for vision–language tasks, remote sensing analysis, and multimodal learning.
The prompt template used for hyperspectral image captioning is as follows:
{mdframed}
[backgroundcolor=gray!5, linecolor=black] You are an AI visual assistant capable of describing a scene based on a segmentation map. The map represents different land cover types using specific colors. The legend is as follows:
- [$color\_i$] corresponds to [$label\_i$], occupying [$percent\_i$]of the image.
Do not mention any colors, color coding, or technical details. Use the given class names. Only mention land cover types in the color legend.
Generate a brief and natural description of the scene by refining: “The hyperspectral image contains [$presented\_labels$] land types.” Provide a concise description of their spatial distributions (e.g., left, right, top, bottom).
To provide a clearer overview of this dataset, we showcase the word cloud in Fig. 12. We also showcase the histogram of the words in the NLCD-hyper-caption dataset in Fig. 13.
A.2.5 SAR-ship-caption and IR-ship-caption
For the SAR-ship-caption and IR-ship-caption subsets, we derive short captions from existing datasets to ensure high-quality textual descriptions. For IR-ship-caption, we utilize samples from the HIT-UAV dataset, specifically the data provided in EarthGPT [66], which is originally designed for visual question answering tasks. To generate descriptions for infrared images, we concatenate the corresponding question and answer pairs, transforming them into concise yet informative captions.
Similarly, for the SAR-ship-caption dataset, we generate captions for SAR images using three key datasets: HRSID [51], SSDD [65], and AIR-SARShip-2.0 [42]. These datasets provide extensive annotated SAR imagery, enabling the creation of meaningful captions. For further details on the captioning methodology, please refer to [66].
A.2.6 ChatEarthNet captions
Figure 14: Word cloud for the ChatEarthNet-caption datasets.
Figure 15: Word histogram for the ChatEarthNet-caption datasets.
Regarding the ChatEarthNet-S2-caption and ChatEarthNet-SAR-caption datasets, we reuse the detailed descriptions provided in the original ChatEarthNet dataset. Specifically, we use Pixtral 12B [2] to summarize these detailed descriptions into short image captions and generate question-answer pairs for visual question answering tasks. The word cloud for ChatEarthNet-caption datasets is presented in Fig. 14. To provide more insights into the dataset, we also show the histogram in Fig. 15. The prompt is as follows: {mdframed}[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
A.3 More dataset examples
These paired samples not only capture scene-level descriptions but also highlight fine-grained properties, such as object positions and fine-grained land cover types. Thereby the dataset enables the development of foundation models proficient in downstream performance for land cover mapping, flood monitoring, and object detection under challenging conditions (e.g., night or cloudy weather).
In Fig. 16 and Fig. 17, we provide more examples to provide a better overview of our dataset.
Figure 16: More examples of the image-text pair of GeoLangBind-2M, our multimodal image-text dataset.
Figure 17: More examples of the image-text pair of GeoLangBind-2M, our multimodal image-text dataset.
A.4 Ablation studies on distillation loss terms
Table 11: Distillation loss balancing weights vs. accuracy
$\alpha_{s}$ (SigLIP) $\alpha_{d}$ (DinoV2) $\alpha_{v}$ (ViT) aid (Top-1) eurosat (Top-1)
0 0 0 59.10 32.96
1.0 0.0 0.0 61.42 33.63
1.0 1.0 0.0 65.36 35.26
1.0 1.0 1.0 68.75 38.63
2.0 0.0 0.0 64.85 33.19
2.0 0.5 0.5 64.70 25.30
2.0 1.0 1.0 69.20 36.67
Tab. 11 presents an ablation study on the impact of different distillation loss balancing weights ($\alpha_{s}$, $\alpha_{d}$, $\alpha_{v}$) on classification accuracy for the AID and EuroSAT datasets. The baseline model (no distillation) achieves 59.10% (AID) and 32.96% (EuroSAT). Adding SigLIP distillation ($\alpha_{s}=1.0$) improves performance, while combining SigLIP and DinoV2 ($\alpha_{s}=1.0,\alpha_{d}=1.0$) further enhances results, especially for AID (65.36%). Introducing ViT distillation ($\alpha_{v}=1.0$) boosts EuroSAT accuracy, peaking at 38.63% when $\alpha_{s}=\alpha_{d}=\alpha_{v}=1.0$. In this work, we adopt the best overall configuration, $\alpha_{s}=2.0,\alpha_{d}=1.0,\alpha_{v}=1.0$, achieves the highest AID accuracy (69.20%) and strong EuroSAT performance (36.67%). This demonstrates the effectiveness of multi-teacher distillation.
A.5 Confusion matrices of zero-shot classification tasks
Figs. 18, 19, 20, 21, 22, 24, 23 and 25 present the confusion matrices of our GeoLangBind-L-384 model on all eight datasets for zero-shot classification. These confusion matrices reveal that common model mistakes largely occur on classes for which humans would also have trouble distinguishing, or for which classes have similar semantic meaning. For instance, the model commonly confuses “sparse residential” and “medium residential” (RESISC45), “annual crop” and “permanent crop” (EuroSAT), “aquaculture land” and “pond” (SkyScript), “chaparral” and “desert” (RESISC45), “mine” and “quarry” (SkyScript, Million-AID), “recreational facility” and “stadium” and “playground” (fMoW, AID), and “ship” and “harbor” and “ferry terminal” (RESISC45, PatternNet).
In Fig. 26, we show the comparison of confusion matrices between SkyScript (ViT-L) and GeoLangBind-L-384 on fine-grained zero-shot classification tasks. It can be clearly seen that our model performs better.
Figure 18: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on SkyScript dataset.
Figure 19: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on AID dataset.
Figure 20: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on EuroSAT dataset.
Figure 21: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on fMoW dataset.
Figure 22: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on Million-AID dataset.
Figure 23: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on PatternNet dataset.
Figure 24: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on the RESISC45 dataset.
Figure 25: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on RSICB dataset.
Figure 26: Comparison of confusion matrices between SkyScript-Large and GeoLangBind-L-384 on the fine-grained zero-shot classification task.
A.6 More visualization of features
To better understand the representations learned by different models, Fig. 27 visualizes the feature maps extracted from RemoteCLIP (ViT-L), SkyScript (ViT-L), and GeoLangBind-L-384 across two datasets: m-chesapeake and m-nz-cattle from GEO-Bench. The m-chesapeake dataset primarily focuses on land cover classification, while the m-nz-cattle dataset captures cattle in remote sensing images. The feature visualization highlights key differences in spatial structure, consistency, and detail preservation across models.
Spatially structured features
GeoLangBind produces features with higher spatial resolution, enabling finer spatial distinctions compared to RemoteCLIP and SkyScript. The feature maps exhibit well-defined structures that align with meaningful spatial regions in the original images, such as roads, buildings, and vegetation.
Consistency across semantic classes
The feature maps from GeoLangBind demonstrate greater consistency across images with similar semantic content. As observed, buildings and roads consistently exhibit high activations, while tree-covered areas show lower activations. This structured activation pattern is apparent across multiple samples, suggesting that GeoLangBind learns more stable and semantically aligned representations.
Preservation of details
GeoLangBind retains finer details within its feature maps, capturing small-scale variations and object boundaries more effectively than the baselines. This advantage is particularly evident in the m-nz-cattle dataset, where cattle are localized with sharper feature activations, as highlighted by the red bounding boxes. RemoteCLIP and SkyScript, on the other hand, exhibit more diffused and less distinct feature responses, making object-level distinctions less precise.
Figure 27: Visualization of features from RemoteCLIP (ViT-L), SkyScript (ViT-L), and GeoLangBind-L-384 on the m-chesapeake and m-nz-cattle datasets from GEO-Bench.

### v2

@@
-GeoLangBind: Unifying Earth Observation with Agglomerative Vision–Language Foundation Models
+DOFA-CLIP: Multimodal Vision–Language Foundation Models for Earth Observation
 Abstract
-Earth observation (EO) data, collected from diverse sensors with varying imaging principles, present significant challenges in creating unified analytical frameworks.
-We present GeoLangBind, a novel agglomerative vision–language foundation model that bridges the gap between heterogeneous EO data modalities using language as a unifying medium.
-Our approach aligns different EO data types into a shared language embedding space, enabling seamless integration and complementary feature learning from diverse sensor data.
-To achieve this, we construct a large-scale multimodal image–text dataset, GeoLangBind-2M, encompassing six data modalities.
-GeoLangBind leverages this dataset to develop a zero-shot foundation model capable of processing arbitrary numbers of EO data channels as input.
-Through our designed Modality-aware Knowledge Agglomeration (MaKA) module and progressive multimodal weight merging strategy, we create a powerful agglomerative foundation model that excels in both zero-shot vision–language comprehension and fine-grained visual understanding.
-Extensive evaluation across 23 datasets covering multiple tasks demonstrates GeoLangBind’s superior performance and versatility in EO applications, offering a robust framework for various environmental monitoring and analysis tasks.
-The dataset and pretrained models will be publicly available.
-1 Introduction
-Figure 1: GeoLangBind leverages language as a bridge to unify heterogeneous remote sensing data, enabling a common representation space across diverse modalities for improved multimodal learning.
-Earth observation (EO) technologies have undergone remarkable advancement in recent years, generating vast amounts of multimodal geospatial data spanning different sensors and scales [54, 55, 70, 59].
-In response to the exponential growth of remote sensing data, researchers have proposed foundation models pretrained on large-scale EO data, including SatMAE [12], Scale-MAE [38], HyperSIGMA [46], SSL4EO [47, 39], and SpectralGPT [20].
-However, these models typically focus on a single EO data modality, limiting their ability to leverage complementary characteristics across different sensor types.
-To address this limitation, recent works such as OFA-Net [58], DOFA [57], AnySat [5], OmniSat [6], and FoMo-Net [8] have introduced foundation models capable of learning from multiple EO modalities within a unified framework.
+Earth observation (EO) spans a broad spectrum of modalities, including optical, radar, multispectral, and hyperspectral data, each capturing distinct environmental signals.
+However, current vision-language models in EO, particularly CLIP-based variants, remain confined to individual modalities, limiting generalization and scalability across diverse tasks.
+We present DOFA-CLIP (Dynamic-One-For-All CLIP), a unified vision-language foundation model that dynamically adapts to EO modalities with flexible spectral configurations through a single Transformer backbone.
+Our approach introduces three key contributions: 1) the construction of GeoLangBind-2M, a large-scale EO image–text dataset covering six heterogeneous modalities with rich natural language descriptions; 2) a novel training strategy called VECT (Vision-models Enhanced Contrastive Text-image pretraining), which enhances the spatial awareness of CLIP features with multiple vision foundation models; and 3) a Modality-aware Knowledge Agglomeration (MaKA) module that refines feature distillation with modality-specific awareness.
+DOFA-CLIP achieves state-of-the-art zero-shot performance across a wide range of EO benchmarks, including unseen modalities and a diverse number of input spectral bands.
+Together, these contributions establish a scalable foundation for multimodal EO understanding and open new avenues for integrating heterogeneous EO data with large language models.
+Code and datasets will be released.
+Introduction
+Figure 1: DOFA-CLIP employs a single unified vision Transformer to align heterogeneous EO data modalities using language as the grounding signal for cross-modal understanding and zero-shot generalization.
+Earth observation (EO) technologies have undergone remarkable advancement in recent years, generating vast amounts of multimodal geospatial data spanning different sensors and scales (Xia et al.
+2018; Zhu et al.
+2017; Xiong et al.
+2024c).
+In response to the exponential growth of remote sensing data, researchers have proposed foundation models (Cong et al.
+2022; Reed et al.
+2023; Wang et al.
+2024a, 2023a; Stewart et al.
+2023; Hong et al.
+2024) pretrained on large-scale EO data for multiple downstream tasks.
+However, these models typically focus on individual EO data modality, limiting their ability to leverage complementary characteristics across different sensor types (Zhu et al.
+2024).
+To address this limitation, recent works such as OFA-Net (Xiong et al.
+2024b), DOFA (Xiong et al.
+2024a), AnySat (Astruc et al.
+2024a), OmniSat (Astruc et al.
+2024b), and FoMo-Net (Bountos, Ouaknine, and Rolnick 2023) have introduced foundation models capable of learning from multiple EO modalities within a unified framework.
 These approaches have demonstrated promising results in multimodal EO representation learning.
 However, their focus remains on unsupervised visual pretraining, lacking a structured way to align EO data with language for better interpretability and retrieval.
-Meanwhile, multimodal large language models like GPT [33], LLaMA [44], LLaVA [29], and DeepSeek [27] have shown remarkable capabilities in handling diverse and complex tasks.
-Vision-language foundation models, such as CLIP [36] and SigLIP [64], have played a crucial role as vision encoders of such multimodal large language models.
-Inspired by these advances, EO-specific adaptations like RemoteCLIP [28], RS5M [67], and SkyScript [50] have attempted to bring contrastive learning to remote sensing.
-However, these models are primarily limited to high-resolution RGB optical imagery, and can not handle other diverse EO modalities (e.g., multispectral data, Synthetic Aperture Radar data) that are critical for many applications.
-Training a single vision–language foundation model capable of supporting multiple heterogeneous EO data modalities presents several challenges: 1) the lack of comprehensive multimodal EO datasets with aligned textual descriptions, 2) the design of a flexible architecture to handle variable-length input channels across different modalities, and 3) addressing the imbalance among different modalities in the training data.
-To address such challenges, we propose GeoLangBind, a unified multimodal vision–language foundation model that handles diverse data modalities and aligns them to the shared language feature space, as illustrated in Fig.
+Meanwhile, multimodal large language models like GPT (OpenAI 2023), LLaVA (Liu et al.
+2023), and Pixtral (Agrawal et al.
+2024) have shown remarkable capabilities in handling diverse and complex tasks.
+Vision-language foundation models, such as CLIP (Radford et al.
+2021) and SigLIP (Zhai et al.
+2023), have played a crucial role as vision encoders of such multimodal large language models.
+Inspired by these advances, EO-specific adaptations like RemoteCLIP (Liu et al.
+2024), RS5M (Zhang et al.
+2024b), RS-CLIP (Li et al.
+2023), and SkyScript (Wang et al.
+2024c) have attempted to bring vision-language contrastive learning to remote sensing.
+However, existing CLIP models for EO have limitations: 1) These models are usually limited to high-resolution RGB imagery and cannot handle modalities like multi/hyper-spectral or SAR, which are vital for many applications.
+2) Compared with vision models like DINOv2 (Oquab et al.
+2023), CLIP models are unsuitable for dense understanding tasks due to the lack of spatial awareness.
+Training a single powerful vision–language model capable of handling heterogeneous EO data modalities raises several core research questions (RQ). These include:
+[…]
+    RQ1: How can we build a large-scale, semantically rich image–text dataset for diverse heterogeneous EO modalities?
+[…]
+    RQ2: How can we efficiently train large CLIP models for EO with better spatial awareness?
+[…]
+    RQ3: How can a single CLIP model handle heterogeneous EO data while preserving modality-specific characteristics?
+To address such challenges, we propose Dynamic-One-For-All CLIP (DOFA-CLIP), a unified multimodal vision–language foundation model that aligns diverse EO modalities to the language space using a single Transformer, as shown in Fig.
 1.
-In addition, GeoLangBind enhances fine-grained semantic understanding via knowledge agglomeration [37] from multiple teacher models beyond contrastive learning.
-To achieve this, we combine the following key designs: 1) a wavelength-aware dynamic encoder to handle multiple data modalities, 2) a modality-aware knowledge agglomeration module, and 3) progressive weight-space merging [53] to scale the model efficiently across multiple modalities.
 Our contributions are as follows:
   1. 1.
-    We construct GeoLangBind-2M, a comprehensive dataset of two million image–text pairs spanning six distinct EO data modalities to facilitate the training of multimodal geospatial foundation models.
+    We introduce GeoLangBind-2M, a large-scale EO dataset comprising two million image–text pairs across six heterogeneous modalities, providing high-quality text descriptions for multimodal vision–language training.
   2. 2.
-    We propose an effective Modality-aware Knowledge Agglomeration (MaKA) module to enhance continuous pre-training and improve fine-grained image understanding.
+    We propose Vision-models Enhanced Contrastive Text-image pretraining (VECT), a simple yet effective training strategy that enhances the fine-grained spatial awareness of CLIP models with open vision foundation models.
   3. 3.
-    We design a progressive weight merging strategy that enables scalable model adaptation across multiple modalities, outperforming traditional data mixing approaches.
+    We develop the Modality-aware Knowledge Agglomeration (MaKA) module to support modality-adaptive representation distillation within a unified Transformer.
   4. 4.
-    We conduct extensive experiments on zero-shot classification, semantic segmentation, and cross-modal image retrieval, demonstrating the superiority of GeoLangBind across diverse tasks.
-Benchmarking results indicate that GeoLangBind achieves state-of-the-art performance while being flexible in handling diverse data modalities with any number of spectral bands.
-Figure 2: Visualization of data samples from our GeoLangBind-2M.
+    We conduct extensive evaluations across zero-shot classification and cross-modal retrieval tasks, demonstrating state-of-the-art performance and better spatial awareness on both seen and unseen EO modalities.
+Figure 2: Visualization of data samples from GeoLangBind-2M.
 The dataset includes imagery from six different sensors and modalities: Sentinel-2 multispectral, Sentinel-1 SAR, EnMAP hyperspectral, elevation maps, infrared imagery, and aerial imagery.
 Each sample is paired with textual descriptions capturing key land cover types, objects, and geographic features.
-2 Related work
+Related work
 Vision foundation models in Earth observation
-Existing pretrained models predominantly focus on RGB data, such as GFM [31], Scale-MAE [38], and Cross-Scale MAE [43].
-Others specialize in multispectral imagery, with FG-MAE [48] and SatMAE [12] tailored for Sentinel-2 data, while SSL4EO-L [39] is trained on Landsat imagery.
-SpectralGPT [20] leverages a 3D generative transformer for spectral data analysis.
-Beyond optical data, CROMA [14] introduces dual unimodal encoders for multispectral and Synthetic Aperture Radar (SAR) data, using a cross-modal radar–optical transformer to learn unified deep representations.
-DeCUR [49] enhances bimodal learning by decoupling unique and shared features across modalities.
-Meanwhile, Satlas [7] assembles a large-scale multi-sensor dataset, pretraining separate models for each sensor.
-DOFA [57] introduces a neural plasticity-inspired hypernetwork that adapts to different sensor wavelengths, enabling joint training on five distinct sensors.
-OmniSat [6] proposes to fuse features across multiple EO modalities to learn robust multimodal representations without labeled data.
-SkySense [18] utilizes separate visual encoders for different data modalities and designs a dedicated module for multimodal fusion.
-Zero-shot foundation models
-Zero-shot foundation models like CLIP [36] have significantly advanced computer vision by enabling models to generalize across tasks without explicit task-specific training.
-Building upon CLIP, SigLIP [64] introduces a pairwise sigmoid loss function to enlarge the training batch and enhance the model performance.
+Existing pretrained models predominantly focus on RGB data, such as GFM (Mendieta et al.
+2023), Scale-MAE (Reed et al.
+2023), and Cross-Scale MAE (Tang et al.
+2024).
+Others specialize in multispectral imagery, with FG-MAE (Wang et al.
+2023b) and SatMAE (Cong et al.
+2022) tailored for Sentinel-2 data, while SSL4EO-L (Stewart et al.
+2023) is trained on Landsat imagery.
+SpectralGPT (Hong et al.
+2024) leverages a 3D generative transformer for spectral data analysis.
+CROMA (Fuller, Millard, and Green 2023) and DeCUR (Wang et al.
+2024b) use self-supervised multimodal learning to learn features across modalities.
+Satlas (Bastani et al.
+2023) assembles a large-scale multi-sensor dataset, pretraining separate models for each sensor.
+DOFA (Xiong et al.
+2024a) introduces a neural plasticity-inspired hypernetwork that adapts to different sensor wavelengths, enabling joint training on five distinct sensors.
+OmniSat (Astruc et al.
+2024b) proposes to fuse features across multiple EO modalities to learn robust multimodal representations without labeled data.
+SkySense (Guo et al.
+2024) utilizes separate visual encoders for different data modalities and designs a dedicated module for multimodal fusion.
+Vision-language foundation models
+Zero-shot foundation models like CLIP (Radford et al.
+2021) have significantly advanced computer vision.
+Building upon CLIP, SigLIP (Zhai et al.
+2023) introduces a pairwise sigmoid loss function to enlarge the training batch and enhance the model performance.
 Zero-shot foundation models have also been investigated in EO applications.
-SkyScript [50] constructs a large-scale image–text dataset by linking remote sensing images with OpenStreetMap [34] semantics.
-RemoteCLIP [28] introduces an image–text dataset by generating captions from existing annotated datasets.
-GeoRSCLIP [67] introduces RS5M, an image–text dataset with 5M pairs, created by filtering existing datasets and generating captions for label-only remote sensing data.
-Despite these advances, existing vision–language models mainly focus on RGB data and have limitations in learning representations from multiple EO data modalities, which are crucial for holistic environmental analysis.
-To unify multiple data modalities, ImageBind [16] aligns six modalities within a shared embedding space.
-LanguageBind [69] extends this by using language as a universal intermediary for modality alignment.
-While these approaches demonstrate the potential of language or vision as unifying mediums, applying them to EO data is still challenging due to the scarcity of multimodal image–text datasets and the diverse scales of Earth systems.
-3 Dataset construction: GeoLangBind-2M
+SkyScript (Wang et al.
+2024c) constructs a large-scale image–text dataset by linking remote sensing images with OpenStreetMap (OpenStreetMap contributors 2017) semantics.
+RemoteCLIP (Liu et al.
+2024) introduces an image–text dataset by generating captions from existing annotated datasets.
+GeoRSCLIP (Zhang et al.
+2024b) introduces RS5M, an image–text dataset with 5M pairs, created by filtering existing datasets and generating captions for label-only remote sensing data.
+Despite these advances, existing vision–language models mainly focus on RGB data and have limitations in learning representations from multiple EO data modalities.
+To unify multiple data modalities, ImageBind (Girdhar et al.
+2023) and TaxaBind (Sastry et al.
+2025) align multiple modalities within a shared embedding space.
+LanguageBind (Zhu et al.
+2023) uses language as a universal intermediary for modality alignment.
+In the era of Large Language Models (LLMs), methods such as GeoChat (Kuckreja et al.
+2024), SkyEyeGPT (Zhan, Xiong, and Yuan 2025), and TeoChat (Irvin et al.
+2024) integrate CLIP-based vision models with LLMs to unify diverse tasks and achieve impressive performance.
+GeoLangBind-2M Dataset
 Table 1: Summary of dataset sources of GeoLangBind-2M.
-* indicates image–text datasets that are created in this work.
+* indicates image–text datasets created in this work.
 SAR = synthetic aperture radar, MSI = multispectral imagery, HSI = hyperspectral imagery, and IR = infrared.
@@
 Total 2,050,983
-GeoLangBind-2M contains a broad range of data modalities including RGB, SAR, multispectral, hyperspectral, infrared, and elevation data, as listed in Tab.
-1.
+In response to RQ1, we propose GeoLangBind-2M, a large-scale image-text dataset covering a broad range of modalities including RGB, SAR, multispectral, hyperspectral, infrared, and elevation data, as listed in Table 1.
 Each image is paired with a high-quality text annotation, as illustrated in Fig.
 2.
-All datasets are implemented using the TorchGeo library [40].
+All datasets are implemented using the TorchGeo library (Stewart et al.
+2022).
 In the following, we introduce the composition of the dataset.
@@
 High-resolution RGB images with textual descriptions form a significant portion of GeoLangBind-2M.
-We incorporate parts of the RemoteCLIP [28] dataset, i.e.
-Seg4 and Det10, which are constructed by generating captions from existing segmentation and detection datasets.
-We also include VRSBench [26], which aggregates multiple object detection datasets.
-SkyScript [50] is a large-scale image–text dataset for EO, from which we include the top 30% highest quality samples after filtering.
-Beyond existing datasets, we introduce Flair2-RGB-caption [15], a newly created image–text dataset.
-Using semantic information as prompts, we employ the Pixtral-12B [2] model to generate captions for detailed land cover features.
+We incorporate parts of the RemoteCLIP (Liu et al.
+2024) dataset: Seg4 and Det10.
+We also include the training split of VRSBench (Li, Ding, and Elhoseiny 2024), which aggregates multiple object detection datasets.
+SkyScript (Wang et al.
+2024c) is a global image–text dataset for EO, from which we include the top 30% highest quality samples.
+In addition to existing datasets, we construct the Flair2-RGB-caption (Garioud et al.
+2023) dataset by leveraging the Pixtral-12B (Agrawal et al.
+2024) model, using semantic information as prompts to generate captions for land cover features.
 SAR image–text datasets
-To enhance the dataset with SAR–text pairs, we create MMFlood-SAR-caption [32], which contains captions emphasizing water extent and flooded regions.
+To enhance the dataset with SAR–text pairs, we create MMFlood-SAR-caption (Montello, Arnaudo, and Rossi 2022), which contains captions emphasizing water extent and flooded regions (more details in supplementary material).
 Additionally, the SAR-ship-caption dataset includes captions describing ships in SAR images.
@@
 Multispectral, hyperspectral, and elevation image–text datasets
-We construct ChatEarthNet-S2-caption by summarizing long, detailed descriptions from the ChatEarthNet dataset [63].
+We construct ChatEarthNet-S2-caption by summarizing long, detailed descriptions from the ChatEarthNet dataset (Yuan et al.
+2024).
 This dataset contains Sentinel-2 multispectral imagery covering more than 10 land cover types.
-Furthermore, we introduce NLCD-hyper-caption, a hyperspectral dataset with 200+ EnMAP bands [17].
+Furthermore, we introduce NLCD-hyper-caption (Braham et al.
+2024), a hyperspectral dataset with 200+ EnMAP bands (Guanter et al.
+2015).
 Like Flair2-RGB-caption, we use Pixtral-12B to generate rich captions, leveraging annotated land cover maps for enhanced descriptiveness.
@@
 Further details on the construction of GeoLangBind-2M are provided in the Supplementary materials.
-Figure 3: GeoLangBind aims to unify and align diverse Earth observation data modalities into the shared language feature space while enhancing fine-grained image understanding.
-The model uses modality-aware knowledge agglomeration (MaKA) to transfer knowledge from multiple teachers.
+Figure 3: DOFA-CLIP aims to align diverse Earth observation data modalities into the language feature space.
+Vision-models Enhanced Contrastive Text-image pretraining (VECT) is proposed to enhance fine-grained image understanding.
+Modality-aware knowledge agglomeration (MaKA) is designed to transfer knowledge from heterogeneous sensors.
 The wavelength-aware dynamic patch embedding module processes multimodal images dynamically.
-4 Methodology
-4.1 GeoLangBind model design
-GeoLangBind aims to unify diverse data modalities and embed them into a shared language feature space while enhancing fine-grained image understanding ability.
-To this end, we design a wavelength-aware dynamic encoder, a modality-aware knowledge agglomeration module, and a progressive weight-space merging approach.
-4.2 Wavelength-aware dynamic encoder
-To handle the diversity of spectral bands across modalities, we adopt the dynamic encoder architecture from DOFA [57], as illustrated in Fig.
+Dynamic-One-For-All CLIP
+Wavelength-aware dynamic encoder
+To handle the diversity of spectral bands across modalities, we adopt the dynamic encoder architecture from DOFA (Xiong et al.
+2024a), as illustrated in Fig.
 3.
@@
 The encoded wavelengths $\mathbf{V}_{\lambda}$ are then transformed using two fully-connected layers with residual connections.
-Next, a lightweight Transformer [45] encoder processes the transformed wavelength embeddings.
+Next, a lightweight Transformer (Vaswani et al.
+2017) encoder processes the transformed wavelength embeddings.
 Specifically, we concatenate the transformed embeddings $\mathbf{V^{\prime}}_{\lambda}$ with the learnable weight query tokens $\mathbf{Q}_{w}$ and a learnable bias query token $\mathbf{Q}_{b}$ as input to the Transformer encoder.
 After the Transformer layer, the output features are passed through two fully-connected layers to generate the dynamic convolutional weights $\mathbf{M}_{w}$ and biases $\mathbf{M}_{b}$.
-Finally, these dynamically generated weights and biases are used for patch embedding (implemented as convolution).
-This process enables the model to handle different data modalities with varying spectral channels using a unified deep encoder.
-4.3 Modality-aware knowledge agglomeration
-We introduce the modality-aware knowledge agglomeration (MaKA) module, which leverages wavelength as a modality-specific condition when distilling features from different data modalities.
-MaKA comprises three key components: a wavelength-aware prompt generator, which encodes wavelength values into a learnable embedding vector; a modality-aware conditional layer normalization, which adapts normalization parameters based on modality-specific characteristics; and a feature projector, which transforms input features to facilitate effective feature matching and knowledge distillation.
-Below, we introduce each of them in detail.
-Wavelength-aware prompt generator
+Finally, the dynamically generated weights and biases are used for patch embedding to handle EO data modalities with varying spectral channels.
+Vision-models enhanced CLIP
+To address RQ2, we propose VECT to enhance the spatial coherence of CLIP features by knowledge distillation from open vision models.
+Specifically, in addition to the contrastive loss, VECT leverages a feature-matching loss to align the intermediate features of DOFA-CLIP with features from multiple pre-trained vision models: SigLIP (Zhai et al.
+2023), DINOv2 (Oquab et al.
+2023), and ViT (Dosovitskiy et al.
+2020).
+As illustrated in Fig.
+3, we extract three RGB channels from multispectral or hyperspectral images as input to the teacher model.
+The student features $\mathbf{F}_{s}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ are then matched to the teacher’s features $\mathbf{F}_{t}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ using a combination of $L_{1}$ loss, cosine embedding loss, and mean squared error (MSE) loss:
+$L_{\text{match}}=L_{L_{1}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{mse}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{cos}}(\mathbf{F}_{s},\mathbf{F}_{t}).$ (1)
+DOFA-CLIP is trained on GeoLangBind-2M using both the pairwise sigmoid loss $L_{\text{siglip}}$ for contrastive learning and distillation loss for feature alignment. The total loss function is:
+$L=L_{\text{siglip}}+\alpha_{s}L_{\text{match}}^{\text{siglip}}+\alpha_{d}L_{\text{match}}^{\text{dinov2}}+\alpha_{v}L_{\text{match}}^{\text{vit}},$ (2)
+where $\alpha_{s},\alpha_{d},$ and $\alpha_{v}$ are balancing factors for feature-matching losses of SigLIP, DINOv2, and ViT, respectively.
+Modality-aware knowledge agglomeration
+Considering RQ3, we introduce the modality-aware knowledge agglomeration (MaKA) module, which is used to translate DOFA-CLIP features in a modality-aware manner to facilitate effective feature matching in VECT.
+MaKA comprises three key components: a wavelength-aware prompt generator, a modality-aware conditional layer normalization, and a feature projector.
+1. Wavelength-aware prompt generator
 Each modality is associated with a set of wavelengths, which we encode using a sine-cosine positional encoding function.
@@
 where $\mathbf{V}_{p}\in\mathbb{R}^{d}$ is the fused modality embedding. This final embedding vector $\mathbf{V}_{p}$ captures modality-specific information derived from the provided wavelengths.
-Modality-aware layer normalization
-Given the feature map from the last layer of the GeoLangBind model, we first interpolate it to match the teacher’s feature resolution, denoted as $\mathbf{F}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$, omitting the batch dimension for simplicity.
+2. Modality-aware layer normalization
+Given the feature map from the last layer of the DOFA-CLIP model, we first interpolate it to match the teacher’s feature resolution, denoted as $\mathbf{F}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$, omitting the batch dimension for simplicity.
 With $\mathbf{F}$ and a modality prompt $\mathbf{V}_{p}\in\mathbb{R}^{d}$, our goal is to perform layer normalization that dynamically adapts to the modality.
-To achieve this, we introduce modality-aware conditional layer normalization, formulated as follows.
+To achieve this, we introduce modality-aware conditional layer normalization as follows.
 First, we derive two learnable modulation vectors, $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{d}$, from the modality prompt via a linear transformation:
@@
 where $\odot$ denotes the Hadamard product. Note that we use the modality-aware layer normalization to learn the residual information specific to each modality.
-Projector module
+3. Projector module
 A convolutional layer is applied to adjust the feature representation $\mathbf{F}^{\prime}$ to channel dimensions $d_{t}$ of the corresponding teachers.
 The final representation is then used to compute the feature-matching loss for model distillation.
-Our design incorporates modality awareness into the feature translation process by encoding wavelength information and adapting the layer normalization parameters accordingly.
-This ensures that the learned features are both aligned and modality-specific, enhancing robust adaptation across different input modalities.
-Model training
-To train the model, we use a feature-matching loss to align the translated spatial features with the teacher’s features.
-As illustrated in Fig.
-3, we extract three RGB channels from multispectral or hyperspectral images as input to the teacher model.
-The student features $\mathbf{F}_{s}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ are then matched to the teacher’s features $\mathbf{F}_{t}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ using a combination of $L_{1}$ loss, cosine embedding loss to maximize the cosine similarity, and mean squared error (MSE) loss:
-$L_{\text{match}}=L_{L_{1}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{mse}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{cos}}(\mathbf{F}_{s},\mathbf{F}_{t}).$ (5)
-We utilize three pre-trained models as teachers: SigLIP [64], DINOv2 [35], and ViT [13].
-During training, the GeoLangBind model is initialized with SigLIP pre-trained weights and trained on GeoLangBind-2M using both the pairwise sigmoid loss $L_{\text{siglip}}$ for contrastive learning and distillation loss for feature matching.
-The total loss function is:
-$L=L_{\text{siglip}}+\alpha_{s}L_{\text{match}}^{\text{siglip}}+\alpha_{d}L_{\text{match}}^{\text{dinov2}}+\alpha_{v}L_{\text{match}}^{\text{vit}},$ (6)
-where $\alpha_{s},\alpha_{d},$ and $\alpha_{v}$ are balancing factors for feature-matching losses of SigLIP, DINOv2, and ViT, respectively.
+This design ensures that the learned features are both aligned and modality-specific, enhancing robust adaptation across different input modalities.
+To address the data imbalance issue, we employ a weight space merging strategy (Ilharco et al.
+2022).
+Since RGB data is the most dominant modality, we split the dataset into two parts: RGB data and non-RGB data.
+We first train two separate DOFA-CLIP models on these subsets.
+Then, we merge their weights to obtain a unified model that can handle both RGB and other modalities.
+Since the original SigLIP model is trained on massive data, we also merge the original SigLIP weights with our model.
+For clarity, let $\theta_{\text{siglip}}$ be the original SigLIP model weights, $\theta_{\text{rgb}}$ be the DOFA-CLIP model weights trained on the RGB subset, $\theta_{\text{others}}$ be the DOFA-CLIP model weights trained on the other five modalities.
+In the first step, we merge $\theta_{\text{siglip}}$ and $\theta_{\text{rgb}}$ using a simple linear weight merging strategy:
+$\theta^{*}=(1-m_{1})\,\theta_{\text{siglip}}+m_{1}\,\theta_{\text{rgb}},$ (7)
+where $m_{1}$ controls the weighting ratio between the two models. Next, we merge the intermediate weights $\theta^{*}$ with $\theta_{\text{others}}$ to obtain the final model:
+$\theta=(1-m_{2})\,\theta^{*}+m_{2}\,\theta_{\text{others}},$ (8)
+where $m_{2}$ determines the contribution of the non-RGB modalities.
+We conduct ablation studies by varying $m_{1}$ and $m_{2}$ to search for optimal weighting ratios.
+While more advanced merging methods exist (Akiba et al.
+2025), we use a linear approach to assess its effectiveness on heterogeneous EO data, leaving sophisticated strategies for future work.
 Table 2: Zero-shot comparison of various models on scene and fine-grained classification tasks. Bold values indicate the best performance.
@@
 SkyCLIP-50 52.98 70.90 33.30 19.24 62.69 72.18 66.67 46.20 53.02 26.00 38.00 67.73
-[…]
+DOFA-CLIP-B-224 70.39 77.60 52.30 20.17 64.91 76.68 67.21 49.53 59.85 44.33 20.00 72.73
 Large CLIP-original 55.06 69.25 41.89 26.19 57.88 71.39 66.70 43.02 53.76 37.50 25.40 42.73
@@
 SkyCLIP-50 70.89 71.70 51.33 27.12 67.45 80.88 70.94 50.09 59.93 46.83 35.80 67.50
-[…]
-Table 3: Zero-shot comparison of various models on GeoBench datasets. Bold values indicate the best performance.
+DOFA-CLIP-L-384 76.83 75.50 59.04 29.10 70.16 80.17 73.15 51.62 64.45 61.83 26.00 81.36
+Table 3: Zero-shot comparison on GeoBench datasets with non-RGB data. Bold values indicate the best performance.
 Model (Base version) m-bigearthnet m-so2sat m-forestnet
@@
 SkyCLIP-50 39.91 19.58 20.32 11.97 - 10.78 -
-[…]
-4.4 Progressive Multimodal Weight Merging
-To address the data imbalance issue, we propose a post-training weight-space merging strategy to scale the model to diverse data modalities effectively.
-Since RGB data is the most dominant modality, we split the dataset into two parts: data with RGB channels and data with other modalities.
-During training, we first train two separate GeoLangBind models on these subsets.
-After training, we merge their weights to obtain a unified model capable of handling both RGB and other modalities.
-Since the original SigLIP model has been trained on massive data, we also merge the weights of GeoLangBind with the original SigLIP weights to enhance the model.
-Thus, we perform two rounds of weight merging to merge these three model weights: 1) merge the original SigLIP weights with the model trained on the RGB subset of GeoLangBind-2M (GeoLangBind-RGB); 2) merge the intermediate model with the model trained on the remaining five data modalities (GeoLangBind-Others).
-For clarity, let $\theta_{\text{siglip}}$ be the original SigLIP model weights, $\theta_{\text{rgb}}$ be the GeoLangBind model weights trained on the RGB subset, $\theta_{\text{others}}$ be the GeoLangBind model weights trained on the other five modalities.
-In the first step, we merge $\theta_{\text{siglip}}$ and $\theta_{\text{rgb}}$ using a simple linear weight merging strategy:
-$\theta^{*}=(1-m_{1})\,\theta_{\text{siglip}}+m_{1}\,\theta_{\text{rgb}},$ (7)
-where $m_{1}$ controls the weighting ratio between the two models. Next, we merge the intermediate weights $\theta^{*}$ with $\theta_{\text{others}}$ to obtain the final GeoLangBind model:
-$\theta=(1-m_{2})\,\theta^{*}+m_{2}\,\theta_{\text{others}},$ (8)
-where $m_{2}$ determines the contribution of the non-RGB modalities.
-We conduct ablation studies in Sec.
-5.3 by varying $m_{1}$ and $m_{2}$ to search for optimal weighting ratios.
-Although more sophisticated weight merging methods exist [3], we adopt a simple linear approach to evaluate its effectiveness for heterogeneous EO data.
-Investigating advanced merging strategies remains a topic for future work.
-5 Experiments
-5.1 Implementation details
-GeoLangBind is trained using 8× NVIDIA A100 40GB GPUs for the base model and 8× NVIDIA A100 80GB GPUs for the large version.
-Our code is implemented based on OpenCLIP [21].
-For the base version, we use a ViT-Base architecture (patch size 16), initialized with WebLI-pretrained weights [9].
+DOFA-CLIP-B (224) 47.70 20.37 23.69 17.95 14.60 13.60 17.02
+Experiments
+Implementation details
+DOFA-CLIP is trained on 8 NVIDIA A100 40GB GPUs for the base model and 8 NVIDIA A100 80GB GPUs for the large version.
+For the base version, we use a ViT-Base architecture (patch size 16) pretrained on WebLI (Chen et al.
+2022).
 The teacher models are ViT-Base-SigLIP (224), DINOv2-ViT-L/14, and ViT-Huge (ImageNet).
-The large version is based on SoViT-400m[4] (patch size 14, input size 384), with teacher models ViT-So400m-SigLIP (384), DINOv2-ViT-L/14, and ViT-Large (ImageNet).
-Training is conducted for 5 epochs on RGB datasets and 20 epochs on other modalities using AdamW (learning rate 5e-4, weight decay 1e-7).
-The batch size is 140 for the base model and 35 for the large model.
-5.2 Performance comparison
-We follow the evaluation protocols of SkyScript [50] and evaluate the models via zero-shot classification tasks.
-Specifically, the following eleven image scene classification datasets are used: AID [54], EuroSAT [19], fMoW [11], Million-AID [30], PatternNet [68], NWPU-RESISC45 [10], RSI-CB256 [25], as well as the fine-grained attribute classification datasets.
-In addition, we further compare these models on some datasets from the GEO-Bench suite [24]: m-bigearthnet [41], m-so2sat [71], and m-forestnet [22].
+The large version is based on SoViT-400m (Alabdulmohsin et al.
+2023) (patch size 14, input size 384), with teacher models ViT-So400m-SigLIP (384), DINOv2-ViT-L/14, and ViT-Large (ImageNet).
+Training is conducted for 20 epochs using AdamW (learning rate 5e-4, weight decay 1e-7).
+Performance comparison
+We follow the evaluation protocols of Wang et al.
+(2024c) and evaluate the models via zero-shot classification tasks.
+Specifically, the following eleven image scene classification datasets are used: AID (Xia et al.
+2017), EuroSAT (Helber et al.
+2019), fMoW (Christie et al.
+2018), Million-AID (Long et al.
+2021), PatternNet (Zhou et al.
+2018), NWPU-RESISC45 (Cheng, Han, and Lu 2017), RSI-CB256 (Li et al.
+2017), as well as the fine-grained attribute classification datasets.
+In addition, we further compare these models on some datasets from the GEO-Bench suite (Lacoste et al.
+2024): m-bigearthnet (Sumbul et al.
+2019), m-so2sat (Zhu et al.
+2019), and m-forestnet (Irvin et al.
+2020).
 Table 4: Recall@1, recall@5, and recall@10 for cross-modal retrieval on three benchmark datasets. Bold values indicate the best performance.
@@
 RSICD, RSITMD, and/or UCM-caption seen in training
-AMFMN [61] 5.39 15.08 23.40 4.90 18.28 31.44 10.63 24.78 41.81 11.51 34.69 54.87 16.67 45.71 68.57 12.86 53.24 79.43
-LW-MCR-u [60] 4.39 13.35 20.29 4.30 18.85 32.34 9.73 26.77 37.61 9.25 34.07 54.03 18.10 47.14 63.81 13.14 50.38 79.52
-GaLR [62] 6.59 19.85 31.04 4.69 19.48 32.13 14.82 31.64 42.48 11.15 36.68 51.68
+AMFMN (Yuan et al. 2022a) 5.39 15.08 23.40 4.90 18.28 31.44 10.63 24.78 41.81 11.51 34.69 54.87 16.67 45.71 68.57 12.86 53.24 79.43
+LW-MCR-u (Yuan et al. 2021) 4.39 13.35 20.29 4.30 18.85 32.34 9.73 26.77 37.61 9.25 34.07 54.03 18.10 47.14 63.81 13.14 50.38 79.52
+GaLR (Yuan et al. 2022b) 6.59 19.85 31.04 4.69 19.48 32.13 14.82 31.64 42.48 11.15 36.68 51.68
 RSICD, RSITMD, and/or UCM-caption not seen in training
@@
 SkyCLIP-30 8.97 24.15 37.97 5.85 20.53 33.53 11.73 33.19 47.35 10.19 32.47 49.08 38.57 84.29 93.81 31.83 64.19 81.96
-[…]
-Table 5: Partial fine-tuning results on six segmentation tasks.
-All models are trained with a frozen backbone for 20 epochs.
-Reported numbers are mean intersection over union (mIoU).
-Missing values are due to the inability of the model to adapt to this domain.
-Method
-Models
-Backbone
-m-pv4ger-seg
-m-nz-cattle
-m-NeonTree
-m-cashew-plant
-m-SA-crop
-m-chesapeake
-Supervised DeepLabv3 ResNet101 93.4 67.6 53.9 48.6 30.4 62.1
-U-Net ResNet101 94.1 80.5 56.6 46.6 29.9 70.8
-Self-supervised Vision FMs rand. init. ViT-B 81.7 74.1 51.7 32.4 29.0 47.1
-Scale-MAE [38] ViT-L 83.5 76.5 51.0 - - 61.0
-GFM [31] Swin-B 92.0 75.0 51.1 - - 63.8
-Cross-Scale MAE [43] ViT-B 83.2 77.9 52.1 - - 52.3
-CROMA [14] ViT-B - - - 30.1 31.4 -
-DOFA [57] ViT-B 94.5 81.4 58.8 51.5 33.0 65.3
-DOFA [57] ViT-L 95.0 81.8 59.4 56.9 32.1 66.3
-Zero-shot FMs RemoteCLIP [28] ViT-L 94.1 79.8 56.2 37.9 31.3 59.5
-Skyscript-20 [50] ViT-L 93.9 79.2 57.2 38.0 30.6 61.5
-GeoLangBind-B-224 ViT-B 94.3 80.7 56.1 53.6 31.8 62.7
-GeoLangBind-L-384 Large 94.7 82.2 59.0 45.8 32.6 63.0
+DOFA-CLIP-L 8.42 25.16 37.05 8.15 24.90 37.73 13.94 30.31 44.47 13.45 38.70 55.78 43.33 81.43 93.81 35.28 71.62 87.80
 Overall zero-shot performance
-As shown in Tab.
-2, our GeoLangBind models establish new state-of-the-art results in zero-shot remote-sensing image classification.
-Compared with existing CLIP-based baselines such as CLIP-original, RemoteCLIP, CLIP-laion-RS, and the various SkyCLIP configurations, GeoLangBind-B-224 (Base) consistently outperforms all counterparts on scene-level tasks.
+As shown in Table 2, DOFA-CLIP establishes new state-of-the-art results in zero-shot remote-sensing image classification.
+Compared with existing CLIP-based baselines such as CLIP-original, RemoteCLIP, CLIP-laion-RS, and the various SkyCLIP configurations, DOFA-CLIP-B-224 consistently outperforms all counterparts.
 Its average accuracy across the eight scene datasets is significantly higher than that of the other base models.
-The Larger variant of GeoLangBind with higher resolution yields further accuracy gains.
-This highlights that both capacity and resolution play vital roles in capturing the spatial and spectral details crucial for complex remote-sensing tasks.
-On three fine-grained roof-attribute classification tasks, both GeoLangBind variants achieve strong accuracy on roof shape and surface, significantly outperforming other methods.
+The Larger variant of DOFA-CLIP with higher resolution yields further accuracy gains.
+On three fine-grained roof-attribute classification tasks, both DOFA-CLIP variants achieve strong accuracy on roof shape and surface, significantly outperforming other methods.
 Although the results on smoothness remain more modest, our approach still provides a clear improvement over existing approaches.
-Zero-shot classification on GEO-Bench
-The zero-shot classification results on three GEO-Bench datasets are presented in Tab.
-3.
-GeoLangBind demonstrates competitive or superior zero-shot multi-label classification performance compared to existing models on the m-bigearthnet dataset (43 classes).
+Zero-shot classification on multispectral imagery
+The zero-shot classification results on three GEO-Bench datasets are presented in Table 3.
+DOFA-CLIP demonstrates competitive or superior zero-shot multi-label classification performance compared to existing models on the m-bigearthnet dataset (43 classes).
 For the m-so2sat dataset (17 classes), when utilizing five spectral bands, the performance decreases slightly due to the significantly less multispectral training data compared to RGB.
 Note that competitive models cannot process 5-band images as input for zero-shot classification.
-On the m-forestnet dataset (12 classes), GeoLangBind substantially outperforms other models and shows remarkable accuracy improvements when using five spectral bands as input.
-Notably, the m-forestnet dataset uses Landsat-8 satellite data, which is not represented in our training dataset.
+On the m-forestnet dataset (12 classes), DOFA-CLIP substantially outperforms other models and shows remarkable accuracy improvements when using five spectral bands as input.
+Notably, the m-forestnet dataset uses Landsat-8 data, which is unseen in our training dataset.
 Cross-modal retrieval performance
-Tab.
-4 compares the GeoLangBind model against recent methods on three benchmark caption datasets: RSICD, RSITMD, and UCM-caption.
+Table 4 compares DOFA-CLIP against recent methods on three benchmark caption datasets: RSICD, RSITMD, and UCM-caption.
 We report Recall@1, Recall@5, and Recall@10 for both image-to-text and text-to-image retrieval.
-Despite the inherent difficulty of remote-sensing cross-modal tasks, GeoLangBind-L consistently achieves competitive or higher recall scores than most competitive baselines, including CLIP-original, CLIP-laion-RS, and SkyCLIP-30.
-Semantic segmentation performance
-Tab.
-5 shows that GeoLangBind models work well across six segmentation tasks despite being zero-shot vision–language models.
-For all models, the backbone remains frozen while UPerNet [56] is trained as the segmentation head.
-GeoLangBind-B-224 achieves strong mIoU scores on m-pv4ger-seg and m-nz-cattle datasets.
-GeoLangBind-L-384 further improves results across all datasets, achieving competitive or better performance than vision foundation models.
-These results demonstrate GeoLangBind’s versatility and ability on fine-grained pixel-level understanding tasks.
+Despite the inherent difficulty of remote-sensing cross-modal tasks, DOFA-CLIP-L consistently achieves competitive or higher recall scores than most competitive baselines, including CLIP-original, CLIP-laion-RS, and SkyCLIP-30.
 Table 5: Ablation experiments of models on the AID zero-shot classification dataset.
@@
 SigLIP 65.95 96.20
-[…]
-SigLIP+GeoLangBind-B-RGB 77.70 95.50
-SigLIP+GeoLangBind-B-RGB 72.00 95.80
-[…]
-SigLIP+GeoLangBind-B-Others 69.50 96.20
-SigLIP+GeoLangBind-B-Others 66.30 91.20
-[…]
-GeoLangBind-B-All-Merging 77.60 96.65
-5.3 Ablation studies
+DOFA-CLIP-B (RGB) 76.90 94.55
+SigLIP+DOFA-CLIP-B (RGB) 77.70 95.50
+SigLIP+DOFA-CLIP-B (RGB) 72.00 95.80
+DOFA-CLIP-B (Others) 64.70 91.95
+SigLIP+DOFA-CLIP-B (Others) 69.50 96.20
+SigLIP+DOFA-CLIP-B (Others) 66.30 91.20
+DOFA-CLIP-B (Mixed) (baseline) 71.60 93.65
+DOFA-CLIP-B (Merging) (ours) 77.60 96.65
+Ablation studies
 Table 6: Ablation experiments of different distillation methods on the segmentation datasets.
@@
 SigLIP 57.2 77.5 52.3
-[…]
-GeoLangBind-B w/ MaKA 62.3 82.2 59.0
-Table 8: Ratio for merging SigLIP and GeoLangBind-B-RGB.
+DOFA-CLIP-B w/o MaKA 61.7 81.5 58.1
+DOFA-CLIP-B w/ MaKA 62.3 82.2 59.0
+Table 7: Ratio for merging SigLIP and DOFA-CLIP-B (RGB).
 $m_{1}$ AID (Top-1) RSI-CB (Top-1) Roof shape (Top-1)
@@
 1.0 76.90 46.72 58.50
-Weight merging
-Tab.
-6 presents the results of our ablation experiments, evaluating the impact of weight merging across different models on the AID zero-shot classification dataset.
-The baseline SigLIP model can be significantly improved by incorporating the GeoLangBind-B-RGB model, trained on the RGB part of GeoLangBind-2M.
-Merging SigLIP with GeoLangBind-B-RGB further enhances performance.
+Weight merging and MaKA
+Table 5 presents the results of our ablation experiments, evaluating the effects of weight merging and MaKA across different models on the AID zero-shot classification dataset.
+The baseline SigLIP model can be significantly improved by incorporating the DOFA-CLIP-B (RGB) model, trained on the RGB part of GeoLangBind-2M.
 The model trained with our MaKA module (for distillation) can largely outperform the non-distilled version.
-This highlights the role of distillation in improving weight fusion.
-GeoLangBind-B-Others, trained on non-RGB modalities performs lower than its RGB counterpart.
+DOFA-CLIP-B (Others), trained on non-RGB modalities, performs lower than its RGB counterpart.
 When merged with SigLIP, the performance can be significantly improved.
-GeoLangBind-B-All is the model trained with all the data modalities mixing together, i.e.
-using data mix instead of weight merging.
-When compared with the GeoLangBind-B-All-Merging, the performance is significantly lower.
-This comparison demonstrates that weight merging is more effective and flexible than direct mixing data modalities.
-Knowledge distillation
-As presented in Tab.
-7, we conducted ablation studies on the MADOS [23], m-nz-cattle [1], and m-NeonTree [52] datasets.
-GeoLangBind-B without MaKA represents a model trained using a baseline distillation method that does not incorporate modality-specific wavelengths as conditioning.
+DOFA-CLIP-B (Mixed) is the model trained with all the data modalities mixed together.
+When compared with the DOFA-CLIP-B (Merging), the performance is significantly lower.
+This demonstrates that weight merging is more effective and flexible than direct mixing of data modalities.
+Figure 4: t-SNE visualization of feature distributions for SigLIP features (left) and DOFA-CLIP-L-384 model (right).
+MaKA enhances segmentation performance
+As presented in Table 6, we conducted ablation studies on the MADOS (Kikaki et al.
+2024), m-nz-cattle (Abuaiadah and Switzer 2022), and m-NeonTree (Weinstein et al.
+2021) datasets.
+DOFA-CLIP-B without MaKA represents a model trained using a baseline distillation method that does not incorporate modality-specific wavelengths as conditioning.
 The results demonstrate the significant effectiveness of the MaKA module, showing consistent performance improvements across all datasets.
+More details are in the supplementary material.
 Weight merging ratios
-In Tab.
-8 and Tab.
-9, a linear search is conducted to determine the optimal ratios for weight merging.
-Based on the results, we select 0.9 for $m_{1}$ to merge the weights of SigLIP and GeoLangBind-RGB.
-We choose 0.5 for merging weights of the intermediate weights and GeoLangBind-Others.
-Thus we set $m_{1}=0.9$ and $m_{2}=0.5$ to derive our final model weights.
-The findings demonstrate that merged weights consistently outperform individual models (when the ratio equals 0 or 1), validating the effectiveness of our weight-merging approach.
-Table 9: Ratio for merging weights of GeoLangBind-B-RGB and GeoLangBind-B-Others.
+In Table 7 and 8, a linear search is conducted to determine the optimal ratios for weight merging.
+Based on the results, we select 0.9 for $m_{1}$ to merge the weights of SigLIP and DOFA-CLIP (RGB).
+We choose 0.5 for merging weights of the intermediate weights and DOFA-CLIP (Others).
+Thus, we set $m_{1}=0.9$ and $m_{2}=0.5$ to derive our final model weights.
+Note that the weight merging is executed a single time using a designated zero-shot classification dataset, and held fixed across all tasks.
+Table 8: Ratio for merging weights of DOFA-CLIP-B (RGB) and DOFA-CLIP-B (Others).
 $m_{2}$ AID (Top-1) RSI-CB (Top-1) Roof shape (Top-1)
@@
 1.0 69.20 34.80 59.67
-Figure 4: t-SNE visualization of feature distributions for the original SigLIP features (left) and GeoLangBind-L-384 model (right).
-Circle markers represent language features, while triangle markers represent image features.
-Figure 5: Visual comparison of deep features from RemoteCLIP, Skyscript, and GeoLangBind model.
-5.4 Feature visualization and analysis
-In the left side of Fig.
+Figure 5: Visual comparison of deep features from RemoteCLIP, Skyscript, and DOFA-CLIP.
+Feature visualization and analysis
+On the left side of Fig.
 4, we present the t-SNE plot of SigLIP features.
-There is a significant separation between the language and image features for hyperspectral, SAR, elevation, and multispectral data modalities.
-This means that SigLIP struggles to align the representations of these data modalities.
-In contrast, features of GeoLangBind-L-384 demonstrate a much stronger feature alignment across all modalities, making it a powerful foundation for Earth observation tasks.
+Compared to the SigLIP features, the features of DOFA-CLIP-L-384 demonstrate a significantly stronger imagery-text alignment across all modalities, making it a powerful foundation for Earth observation tasks.
 In Fig.
-5, we visually compare the feature heatmaps of RemoteCLIP, SkyScript, and GeoLangBind.
-We can see that GeoLangBind features are more spatially structured and preserve more object details.
-6 Conclusion
-We introduce GeoLangBind, a unified vision–language foundation model that bridges heterogeneous EO modalities through a shared language space.
+5, we visually compare the feature heatmaps of RemoteCLIP, SkyScript, and DOFA-CLIP.
+DOFA-CLIP features are more spatially structured and preserve more object details, demonstrating the effectiveness of VECT.
+Conclusion
+We introduce DOFA-CLIP, a unified vision–language foundation model that bridges heterogeneous EO modalities through a shared language space.
 To achieve this, we construct GeoLangBind-2M, a large-scale image–text dataset containing six EO data modalities.
-GeoLangBind consists of a wavelength-aware encoder and a modality-aware knowledge agglomeration module to enhance fine-grained image understanding.
+DOFA-CLIP consists of a wavelength-aware encoder and a modality-aware knowledge agglomeration module to enhance fine-grained image understanding.
 Additionally, progressive weight merging is proposed to scale the model training to multiple data modalities.
-Extensive experiments demonstrate the state-of-the-art performance of GeoLangBind across zero-shot classification, semantic segmentation, and cross-modal retrieval tasks.
-Appendix A Supplementary material
+Extensive experiments demonstrate the state-of-the-art performance of DOFA-CLIP across zero-shot classification, semantic segmentation, and retrieval tasks.
+Supplementary material
 The supplementary material contains the following sections:
@@
     More visualization of the features;
-A.1 Details of the evaluation datasets
-We evaluate GeoLangBind on both datasets that have been established in previous works for direct comparison, as well as additional datasets that highlight the wide-ranging capabilities of the proposed model.
-The following datasets are part of the SkyScript [50] evaluation framework, which we follow for direct comparison.
-  • SkyScript classification dataset [50]: contains 7,000 aerial RGB images with 70 classes and different objects than the SkyScript pretraining dataset.
-  • AID [54]: contains 2,000 aerial RGB images and 30 image scene classes.
-  • EuroSAT [19]: contains 27,000 Sentinel-2 images with 13 spectral bands and 10 image scene classes. We use the dedicated test split that contains 2,700 images.
-  • fMoW [11]: contains 106,081 RGB aerial images with 62 image scene classes.
-  • Million-AID [30]: contains 10,000 RGB aerial images of varying GSD with 51 image scene classes.
-  • PatternNet [68]: contains 30,400 high-resolution RGB aerial images with 6–50 cm GSD and 38 image scene classes.
-  • NWPU-RESISC45 [10]: contains 31,500 aerial RGB images with 0.2–30 m GSD and 45 image scene classes.
-  • RSI-CB [25]: contains 24,747 aerial RGB images and 35 image scene classes.
+Details of the evaluation datasets
+We evaluate DOFA-CLIP on both datasets that have been established in previous works for direct comparison, as well as additional datasets that highlight the wide-ranging capabilities of the proposed model.
+The following datasets are part of the SkyScript (Wang et al. 2024c) evaluation framework, which we follow for direct comparison.
+  • SkyScript classification dataset (Wang et al. 2024c): contains 7,000 aerial RGB images with 70 classes and different objects than the SkyScript pretraining dataset.
+  • AID (Xia et al. 2017): contains 2,000 aerial RGB images and 30 image scene classes.
+  • EuroSAT (Helber et al. 2019): contains 27,000 Sentinel-2 images with 13 spectral bands and 10 image scene classes. We use the dedicated test split that contains 2,700 images.
+  • fMoW (Christie et al. 2018): contains 106,081 RGB aerial images with 62 image scene classes.
+  • Million-AID (Long et al. 2021): contains 10,000 RGB aerial images of varying GSD with 51 image scene classes.
+  • PatternNet (Zhou et al. 2018): contains 30,400 high-resolution RGB aerial images with 6–50 cm GSD and 38 image scene classes.
+  • NWPU-RESISC45 (Cheng, Han, and Lu 2017): contains 31,500 aerial RGB images with 0.2–30 m GSD and 45 image scene classes.
+  • RSI-CB (Li et al. 2017): contains 24,747 aerial RGB images and 35 image scene classes.
 Table 9: GEO-Bench dataset description.
@@
 m-forestnet 332 x 332 12 6464 989 993 6 Landsat-8
-Segmentation m-pv4ger-seg 320 x 320 2 3000 403 403 3 RGB
-m-chesapeake-landcover 256 x 256 7 3000 1000 1000 4 RGBN
-m-cashew-plantation 256 x 256 7 1350 400 50 13 Sentinel-2
-m-SA-crop-type 256 x 256 10 3000 1000 1000 13 Sentinel-2
-m-nz-cattle 500 x 500 2 524 66 65 3 RGB
+Segmentation m-nz-cattle 500 x 500 2 524 66 65 3 RGB
 m-NeonTree 400 x 400 2 270 94 93 5 RGB + Hyperspectral + Elevation
-Additionally, we choose the GEO-Bench [24] suite of datasets that cover a range of relevant remote sensing tasks across different domains, sensors, and geospatial locations.
-Tab.
-10 contains a detailed overview of the number of samples, sensors, and target classes in the GEO-Bench datasets.
+Additionally, we choose the GEO-Bench (Lacoste et al.
+2024) suite of datasets that cover a range of relevant remote sensing tasks across different domains, sensors, and geospatial locations.
+Table 9 contains a detailed overview of the number of samples, sensors, and target classes in the GEO-Bench datasets.
 Figure 6: Word cloud for the Flair2-RGB-caption dataset.
 Figure 7: Word histogram for the Flair2-RGB-caption dataset.
-In addition, we conduct ablation experiments on the Marine Debris and Oil Spill (MADOS) [23] dataset.
+In addition, we conduct ablation experiments on the Marine Debris and Oil Spill (MADOS) (Kikaki et al.
+2024) dataset.
 The MADOS dataset is a high-resolution multispectral Sentinel-2 dataset for marine pollution detection.
@@
 Unlike existing datasets, MADOS enables scalable, generalizable deep learning models for holistic marine pollution monitoring.
-A.2 Details on constructing GeoLangBind-2M
+Details on constructing GeoLangBind-2M
 We construct our multimodal image–text dataset GeoLangBind-2M by assembling existing datasets and curating 8 new datasets. Details on the newly curated sub-datasets are as follows.
-A.2.1 Flair2-RGB-caption
-For the Flair2-RGB-caption dataset, originally designed for semantic segmentation land cover classification, we employed an approach similar to the ChatEarthNet [63] dataset construction.
-We select the training set of 61,711 images from FLAIR #2 [15] and derive contextual information from their corresponding pixel-level segmentation masks.
-This extracted semantic content, paired with the segmentation maps as visual context, serves as the input for generating comprehensive descriptions using the Pixtral 12B [2] language model.
+Flair2-RGB-caption
+For the Flair2-RGB-caption dataset, originally designed for semantic segmentation land cover classification, we employed an approach similar to the ChatEarthNet (Yuan et al.
+2024) dataset construction.
+We select the training set of 61,711 images from FLAIR #2 (Garioud et al.
+2023) and derive contextual information from their corresponding pixel-level segmentation masks.
+This extracted semantic content, paired with the segmentation maps as visual context, serves as the input for generating comprehensive descriptions using the Pixtral 12B (Agrawal et al.
+2024) language model.
 Our semantic extraction process involves calculating the distribution percentages of various terrain categories within each image and formulating these statistics into structured contextual cues.
@@
 In Fig. 6, we illustrate the word cloud for this dataset.
-A.2.2 Flair2-Elevation-caption
+Flair2-Elevation-caption
 The Flair2-Elevation-Caption dataset is constructed to generate detailed textual descriptions of terrain characteristics by leveraging both semantic segmentation maps and elevation maps.
@@
 9.
-A.2.3 MMflood-SAR-caption
+MMflood-SAR-caption
 The MMflood-SAR-caption dataset contains textual descriptions of flood maps generated using Pixtral 12B.
@@
 [backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
-A.2.4 NLCD-hyper-caption
+NLCD-hyper-caption
 The caption generation process for hyperspectral images follows a methodology similar to the FLAIR #2 dataset, ensuring consistency in semantic information extraction and description generation.
@@
 To provide a clearer overview of this dataset, we showcase the word cloud in Fig. 12. We also showcase the histogram of the words in the NLCD-hyper-caption dataset in Fig. 13.
-A.2.5 SAR-ship-caption and IR-ship-caption
+SAR-ship-caption and IR-ship-caption
 For the SAR-ship-caption and IR-ship-caption subsets, we derive short captions from existing datasets to ensure high-quality textual descriptions.
-For IR-ship-caption, we utilize samples from the HIT-UAV dataset, specifically the data provided in EarthGPT [66], which is originally designed for visual question answering tasks.
+For IR-ship-caption, we utilize samples from the HIT-UAV dataset, specifically the data provided in Zhang et al.
+(2024a), which is originally designed for visual question answering tasks.
 To generate descriptions for infrared images, we concatenate the corresponding question and answer pairs, transforming them into concise yet informative captions.
-Similarly, for the SAR-ship-caption dataset, we generate captions for SAR images using three key datasets: HRSID [51], SSDD [65], and AIR-SARShip-2.0 [42].
+Similarly, for the SAR-ship-caption dataset, we generate captions for SAR images using three key datasets: HRSID (Wei et al.
+2020), SSDD (Zhang et al.
+2021), and AIR-SARShip-2.0 (Sun, Sun, and Wang 2020).
 These datasets provide extensive annotated SAR imagery, enabling the creation of meaningful captions.
-For further details on the captioning methodology, please refer to [66].
-A.2.6 ChatEarthNet captions
+For further details on the captioning methodology, please refer to Zhang et al.
+(2024a).
+ChatEarthNet captions
 Figure 14: Word cloud for the ChatEarthNet-caption datasets.
@@
 Regarding the ChatEarthNet-S2-caption and ChatEarthNet-SAR-caption datasets, we reuse the detailed descriptions provided in the original ChatEarthNet dataset.
-Specifically, we use Pixtral 12B [2] to summarize these detailed descriptions into short image captions and generate question-answer pairs for visual question answering tasks.
+Specifically, we use Pixtral 12B (Agrawal et al.
+2024) to summarize these detailed descriptions into short image captions and generate question-answer pairs for visual question answering tasks.
 The word cloud for ChatEarthNet-caption datasets is presented in Fig.
@@
 The prompt is as follows: {mdframed}[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
-A.3 More dataset examples
+More dataset examples
 These paired samples not only capture scene-level descriptions but also highlight fine-grained properties, such as object positions and fine-grained land cover types.
@@
 Figure 17: More examples of the image-text pair of GeoLangBind-2M, our multimodal image-text dataset.
-A.4 Ablation studies on distillation loss terms
+Ablation studies on distillation loss terms
 Table 10: Distillation loss balancing weights vs. accuracy
@@
 2.0 1.0 1.0 69.20 36.67
-Tab.
-11 presents an ablation study on the impact of different distillation loss balancing weights ($\alpha_{s}$, $\alpha_{d}$, $\alpha_{v}$) on classification accuracy for the AID and EuroSAT datasets.
+Table 10 presents an ablation study on the impact of different distillation loss balancing weights ($\alpha_{s}$, $\alpha_{d}$, $\alpha_{v}$) on classification accuracy for the AID and EuroSAT datasets.
 The baseline model (no distillation) achieves 59.10% (AID) and 32.96% (EuroSAT).
@@
 This demonstrates the effectiveness of multi-teacher distillation.
-A.5 Confusion matrices of zero-shot classification tasks
-Figs.
-18, 19, 20, 21, 22, 24, 23 and 25 present the confusion matrices of our GeoLangBind-L-384 model on all eight datasets for zero-shot classification.
-These confusion matrices reveal that common model mistakes largely occur on classes for which humans would also have trouble distinguishing, or for which classes have similar semantic meaning.
+Confusion matrices of zero-shot classification tasks
+Figure 18 to Fig.
+25 present the confusion matrices of our DOFA-CLIP-L-384 model on all eight datasets for zero-shot classification.
+These confusion matrices reveal that common model mistakes largely occur for classes that humans would also have trouble distinguishing, or for classes with similar semantic meanings.
 For instance, the model commonly confuses “sparse residential” and “medium residential” (RESISC45), “annual crop” and “permanent crop” (EuroSAT), “aquaculture land” and “pond” (SkyScript), “chaparral” and “desert” (RESISC45), “mine” and “quarry” (SkyScript, Million-AID), “recreational facility” and “stadium” and “playground” (fMoW, AID), and “ship” and “harbor” and “ferry terminal” (RESISC45, PatternNet).
 In Fig.
-26, we show the comparison of confusion matrices between SkyScript (ViT-L) and GeoLangBind-L-384 on fine-grained zero-shot classification tasks.
+26, we show the comparison of confusion matrices between SkyScript (ViT-L) and DOFA-CLIP-L-384 on fine-grained zero-shot classification tasks.
 It can be clearly seen that our model performs better.
-Figure 18: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on SkyScript dataset.
-Figure 19: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on AID dataset.
-Figure 20: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on EuroSAT dataset.
-Figure 21: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on fMoW dataset.
-Figure 22: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on Million-AID dataset.
-Figure 23: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on PatternNet dataset.
-Figure 24: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on the RESISC45 dataset.
-Figure 25: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on RSICB dataset.
-Figure 26: Comparison of confusion matrices between SkyScript-Large and GeoLangBind-L-384 on the fine-grained zero-shot classification task.
-A.6 More visualization of features
+Figure 18: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on SkyScript dataset.
+Figure 19: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on AID dataset.
+Figure 20: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on EuroSAT dataset.
+Figure 21: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on fMoW dataset.
+Figure 22: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on Million-AID dataset.
+Figure 23: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on PatternNet dataset.
+Figure 24: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on the RESISC45 dataset.
+Figure 25: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on RSICB dataset.
+Figure 26: Comparison of confusion matrices between SkyScript-Large and DOFA-CLIP-L-384 on the fine-grained zero-shot classification task.
+More visualization of features
 To better understand the representations learned by different models, Fig.
-27 visualizes the feature maps extracted from RemoteCLIP (ViT-L), SkyScript (ViT-L), and GeoLangBind-L-384 across two datasets: m-chesapeake and m-nz-cattle from GEO-Bench.
+27 visualizes the feature maps extracted from RemoteCLIP (ViT-L), SkyScript (ViT-L), and DOFA-CLIP-L-384 across two datasets: m-chesapeake and m-nz-cattle from GEO-Bench.
 The m-chesapeake dataset primarily focuses on land cover classification, while the m-nz-cattle dataset captures cattle in remote sensing images.
@@
 Spatially structured features
-GeoLangBind produces features with higher spatial resolution, enabling finer spatial distinctions compared to RemoteCLIP and SkyScript.
+DOFA-CLIP produces features with higher spatial resolution, enabling finer spatial distinctions compared to RemoteCLIP and SkyScript.
 The feature maps exhibit well-defined structures that align with meaningful spatial regions in the original images, such as roads, buildings, and vegetation.
 Consistency across semantic classes
-The feature maps from GeoLangBind demonstrate greater consistency across images with similar semantic content.
+The feature maps from DOFA-CLIP demonstrate greater consistency across images with similar semantic content.
 As observed, buildings and roads consistently exhibit high activations, while tree-covered areas show lower activations.
-This structured activation pattern is apparent across multiple samples, suggesting that GeoLangBind learns more stable and semantically aligned representations.
+This structured activation pattern is apparent across multiple samples, suggesting that DOFA-CLIP learns more stable and semantically aligned representations.
 Preservation of details
-GeoLangBind retains finer details within its feature maps, capturing small-scale variations and object boundaries more effectively than the baselines.
+DOFA-CLIP retains finer details within its feature maps, capturing small-scale variations and object boundaries more effectively than the baselines.
 This advantage is particularly evident in the m-nz-cattle dataset, where cattle are localized with sharper feature activations, as highlighted by the red bounding boxes.
 RemoteCLIP and SkyScript, on the other hand, exhibit more diffused and less distinct feature responses, making object-level distinctions less precise.
-Figure 27: Visualization of features from RemoteCLIP (ViT-L), SkyScript (ViT-L), and GeoLangBind-L-384 on the m-chesapeake and m-nz-cattle datasets from GEO-Bench.
+Figure 27: Visualization of features from RemoteCLIP (ViT-L), SkyScript (ViT-L), and DOFA-CLIP-L-384 on the m-chesapeake and m-nz-cattle datasets from GEO-Bench.